### v2

# Fine-grained Hallucination Detection and Editing for Language Models

Abstract
Large language models (LMs) are prone to generate diverse factually incorrect statements, which are widely called hallucinations. Current approaches predominantly focus on coarse-grained automatic hallucination detection or editing, overlooking nuanced error levels. In this paper, we propose a novel task—automatic fine-grained hallucination detection—and present a comprehensive taxonomy encompassing six hierarchically defined types of hallucination. To facilitate evaluation, we introduce a new benchmark that includes fine-grained human judgments on two LM outputs across various domains. Our analysis reveals that ChatGPT and Llama2-Chat exhibit hallucinations in 60% and 75% of their outputs, respectively, and a majority of these hallucinations fall into categories that have been underexplored. As an initial step to address this, we train Fava, a retrieval-augmented LM by carefully designing synthetic data generations to detect and correct fine-grained hallucinations. On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT on fine-grained hallucination detection by a large margin though a large room for future improvement still exists. Fava’s suggested edits also improve the factuality of LM-generated text, resulting in 5-10% FActScore improvements. [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
1 Introduction
Large language models (LMs; Brown et al. 2020) can generate highly fluent and plausible text. However, these models are prone to produce hallucinations—factually incorrect generations—which often impedes their deployment in real-world applications Mallen et al. (2022). Prior work on hallucinations in natural language generation often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al. (2023). On the contrary, escalating apprehensions have been articulated with respect to LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al. (2023); Huang et al. (2022).
Figure 1: Overview of fine-grained detection system and the proposed taxonomy for hallucinations.
Such prevalence of hallucinations in LM outputs has motivated the development of automatic hallucination detection Min et al. (2023) or editing systems Gao et al. (2022); Chen et al. (2023). However, those existing systems tend to treat diverse types of factual errors uniformly, often simplifying them into binary categories such as “factual” or “not factual” with a focus on certain domains or specific types of errors. While entity-level contradictions (Figure 1 top) are usually evident and can be easily rectified with a single reference, LMs often fabricate entities that lack real-world existence (Figure 1 bottom). This variability underscores the need for a more fine-grained approach to hallucination detection for both model development and human verification, and we argue that a detection system should provide precise identification of hallucination sequences, discern different error types, and suggest potential refinements.
In this work, we introduce a new task, benchmark, and model to facilitate automatic fine-grained hallucination detection (Figure 1). We first introduce a new taxonomy of hallucinations and tasks (Section 3), and collect a high-quality benchmark to test models’ ability to provide fine-grained hallucination detection (Section 4). We then introduce a new model that is capable of providing span-level hallucination detection and editing (Section 5) as the first step towards building reliable fine-grained hallucination detection systems.
Our taxonomy (Figure 2; Table 1) hierarchically classifies factual errors in LM generations into six categories. We specifically focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters. Based on this taxonomy, we propose a new task of fine-grained hallucination detection, in which a system needs to identify locations of factual errors in a response along with fine-grained hallucination types.
We carefully annotate two widely-used LMs’ responses to diverse instruction-following and knowledge-intensive queries and construct the first fine-grained hallucination annotation benchmark. Our benchmark consists of about 400 responses of ChatGPT and Llama2-Chat 70B with span-level annotations of hallucination types based on our proposed taxonomy. Our benchmark reveals on average 1.9 and 3.4 hallucinations per response in ChatGPT and Llama2-Chat, respectively. It also shows that in addition to the widely-studied entity-level errors, other error types like unverifiable sentences or invented concepts make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
Due to the expensive costs of annotations, creating training data for this new task is challenging and our preliminary experiments find proprietary LMs such as ChatGPT perform poorly in fine-grained detection. To overcome those challenges, we design an LM-based synthetic data generation process, resulting in 35k training instances. On this training data, we fine-tune a new retrieval-augmented LM, Fava, which can identify and mark hallucinations at span level using a unified syntax. We compare Fava with state-of-the-art LMs including ChatGPT on fine-grained hallucination detection and editing in LM outputs. Fava significantly outperforms ChatGPT with and without external knowledge by 22.1% on fine-grained hallucination detection and 14.2% on binary detection. Furthermore, Fava effectively fixes factual errors in diverse LM outputs—Fava editing improves FActScore Min et al. (2023) of Alapaca 7, 13B, and ChatGPT output by 4.4%, 9.3% and 3.3%, respectively.
2 Related Work
Hallucinations in natural language generation.
Taxonomy of hallucinations in certain downstream tasks such as summarization Pagnoni et al. (2021), text simplifications Devaraj et al. (2022), or knowledge-grounded dialogues Dziri et al. (2022) have been studied. Such prior taxonomies often assume the existence of specific source text and focus on faithfulness to the source Ji et al. (2023), which is qualitatively different from LMs’ factual hallucinations grounded in world knowledge. In addition, prior taxonomies include many task-specific categories that cannot be directly adapted to describe hallucinations in LM generations.
Detecting and editing hallucinations in LMs.
Several recent works study methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al. (2023); Min et al. (2023), or focusing on entity-level factual errors editing Gao et al. (2022); Chen et al. (2023). Recent surveys have attempted to categorize hallucinations in LM generations, but such categorizations often stretch the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023). In this work, we propose a fine-grained taxonomy for fact-conflicting hallucinations in long-form text generation to improve factual consistency. While prior work often develops factuality verification systems on top of proprietary LMs Chern et al. (2023), we introduce a carefully designed training pipeline for a smaller yet more competitive fine-grained detection system. Unlike prior qualitative analysis on hallucinations in LM responses on certain knowledge-intensive queries (e.g., open-domain QA, biography generation; Liu et al., 2023b; Min et al., 2023), we annotate responses to diverse queries in multiple domains and reveal hallucination distributions vary across domains.
Figure 2: An overview of our fine-grained hallucination taxonomy.
Type Example ChatGPT Llama2
 Entity Lionel Andrés Messi was born on June 12 24, 1987. 46.9% 46.5%
 Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 2.8% 2.5%
 Contradictory Messi has yet to gain captaincy for the Argentina national football team. 12.0% 14.1%
 Invented Messi is also famous for his discovery of the famous airplane kick technique. 4.6% 23.0%
 Subjective Lionel Messi is the best soccer player in the world. 15.0% 8.8%
 Unverifiable In his free time, Messi enjoys singing songs for his family. 18.7% 5.1%
Table 1: Examples of hallucinations with their percentage in our human-annotated data. We also show the percentage of those errors in ChatGPT and LLama2-Chat-70B outputs.
Fact verification.
While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021). Thorne et al. (2018) introduce a large scale dataset based on Wikipedia documents for the training and evaluation of fact verification systems. Other datasets in information-dense domains like fact-checking against scientific articles (Wadden et al., 2020) and news documents (Wang, 2017) have been proposed. Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text. While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims as well.
3 Fine-grained Hallucination Detection
3.1 Focus and Definitions
In this work, we focus on open-ended text generation given information-seeking queries which often require factual knowledge. We define hallucinations as factually incorrect or unverifiable statements given external world knowledge. We operationalize world knowledge as documents from the web that are most relevant to the given statement/query according to a search algorithm. [^2] Errors in common sense, numerical, or logical reasoning are out of the scope of this work.
3.2 Hallucination Taxonomy
Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types. To develop a hallucination taxonomy for open-ended LM generations, we draw inspiration from error categories in prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022) and introduce new categories to describe more complex errors surfacing in LM generations. We conducted a pilot annotation with 9 NLP experts to discuss and refine our taxonomy to ensure good coverage across diverse error types.
Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination. We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
(1a)  Entity : Contradictory entity errors are a sub-category within contradictory statement errors where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
(1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship described in a statement is incorrect. While (1a) often involves errors in nouns, (1b) involves verbs, prepositions, or adjectives, and can be fixed by correcting the incorrect relation.
(1c)  Contradictory : Contradictory statement errors refer to statements that entirely contradict relevant evidence from the web. These are cases where the full sentence is refuted by the information provided in a given reference.
(2)  Invented : Invented errors refer to statements with concepts that do not exist based on world knowledge. These are cases when the LM generates a non-existent or entirely fabricated entity that doesn’t occur in any relevant evidence. This does not include fictional characters in creative work such as books or movies.
(3)  Subjective : Subjective errors refer to an expression or proposition that lacks universal validity and is often influenced by personal beliefs, feelings, opinions, or biases and hence cannot be judged as factually correct. Specifically, this represents cases where the LM generates a statement without any factual proposition grounded in relevant evidence.
(4)  Unverifiable : Unverifiable errors refer to statements that contain factual propositions but cannot be grounded in world evidence. These are cases where the LM generates statements with facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal or private matters).
While  Entity  or  Relation  are often phrase-level and can be fixed by minimal editing erroneous phrases, other error types can be an entire sentence or part of a sentence and should be removed from a response to make it factual.
3.3 Tasks and Metrics
We introduce two new tasks of identifying and editing fine-grained factual errors in LM outputs. Given an input query $x$ and a corresponding LM output $y$, our tasks require systems to identify all of the factual errors. Each error $e$ consists of $(e^{text},e^{type})$, indicating the factually incorrect text spans and their error types among our taxonomies, respectively. We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1) and editing factual errors $e^{text}$ accordingly (Task 2).
Figure 3: Hallucination type distribution of human annotations across 3 datasets and two models.
Task 1: Fine-grained hallucination detection.
In the first fine-grained error detection task, the system is expected to identify fine-grained errors in an LM output. Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$. Given an output $y$ consisting of $L$ sentences, we assume the availability of gold error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE). For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==\texttt{TRUE}]}$ (1)
$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}==\texttt{TRUE}]}$ (2)
Precision indicates how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model. For final score, we compute the F1 scores averaged over six error types as follows:
$\frac{1}{|\mathcal{E}|}\sum_{t\in\mathcal{E}}\frac{2\times\text{Prec}^{t}\times\text{Recall}^{t}}{\text{Prec}^{t}+\text{Recall}^{t}}$ (3)
Fine-grained error detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not. This is similar to FActScore Min et al. (2023), where we predict and evaluate factuality at the sentence level without extracting atomic facts.
Task 2: Hallucination editing.
As shown in Table 1, some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable. In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output. We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore or QAFactEval Fabbri et al. (2022). as follows:$f(\hat{y})-f(y)$, where $f$ indicates an evaluation function for factuality while $f(\hat{y})$ indicates edited output.
4 Benchmark
4.1 Human Annotation on Hallucinations
To facilitate model development in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are, we carefully annotate the first fine-grained hallucination detection benchmark. Our benchmark consists of multi-way fine-grained annotations on two LM responses to queries in multiple domains.
Source prompts.
To evaluate LMs’ factuality in the responses to diverse queries, we collect 150 information-seeking queries, spanning three different categories. First, we sample 50 knowledge-intensive queries from the Open Assistant Köpf et al. (2023) dataset, by prompting GPT-4 to judge whether each query from the dataset requires world knowledge or not. We, the authors further compose 50 instruction-following queries that require more complex reasoning and knowledge. Lastly, to enable more controlled evaluations and analysis, we synthetically create 50 queries that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template. See examples in each category in Appendix Table 6.
Annotation details.
We prompt ChatGPT Ouyang et al. (2022) and Llama2-Chat 70B Touvron et al. (2023) in a zero-shot manner and collect 300 responses to our diverse information-seeking queries. We recruited 15 students (five undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy. Due to the complexity of the task, we conducted 45-minute in-person and virtual training sessions with them. Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation. For each LM response, we assign two annotators. We present model-generated responses, input information-seeking queries, and the top 10 web search documents returned by Google Search given the original input queries. We ask human annotators to first mark a span, annotate a category of error, and suggest an edit to fix the errors for the  Entity  and  Relation  types. Our annotation interface and more details of annotations are available in Appendix A.
Analysis on annotated data.
Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains. 60.8% and 74.7% of the responses of ChatGPT and Llama2-Chat include at least one factual hallucination, respectively.  Entity  is the most widely recognized error type, making up 46.3% of the detected errors. Yet, there are diverse types of errors prevalent like invented statements or contradictory statements which make up 16.8% and 15.4% of the errors detected respectively. We found that the error distributions vary across different source prompts and LMs.  Invented  are more common in Llama2-Chat 70B. We observe fewer errors in Open Assistant subsets, possibly because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How tall is the Burj Khalifa?), which is often memorized by powerful LMs Mallen et al. (2022). On the other hand, the newly annotated instructions and WebNLG data show a higher percentage of errors.
Figure 4: Overview of data creation process in Fava.
Our annotators also agreed that our taxonomy covers most of the error types in LM responses. Furthermore, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in error detection across passages, with 75.1% agreement in error detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages. Appendix Table 8 shows examples of our annotations.
5 Model: Fava
Our benchmark analysis reveals the urgent need to study and detect such diverse types of hallucinations.
We introduce a new retrieval-augmented LM, Fava (FAct Vericaton with Augmentation). Fava is trained on high-quality synthetic training data, and at inference, it identifies and fixes fine-grained hallucinations, incorporating retrieved knowledge.
5.1 Overview
Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$. $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$. Subsequently, the editing model detects and if possible edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$. $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1.
While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2). Reliance on black-box proprietary API models also hurt reproducible. Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.2) and fine-tune a powerful expert LM consisting of 7 billion parameters (Section 5.3).
5.2 Training Data Creation
To train our $\mathcal{M}_{edit}$ model, we require a large number of erroneous LM outputs $y$ paired with edited versions $y^{*}$ with fine-grained error types and edits. Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al. (2022); Wang et al. (2023b); Asai et al. (2023), we introduce a new data creation method grounded on our fine-grained hallucination taxonomies. We prompt GPT-4 Achiam et al. (2023) and ChatGPT to diversify and noise a Wikipedia passage by inserting different types of hallucinations.
In particular, our data creation pipeline consists of three steps: seed passage generations, error insertions, and postprocessing. The first two processes create an input sequence with perturbed diverse error annotations, and postprocessing remaps those special tokens to generate the erroneous output $y$ and output with fine-grained detection results $y^{*}$. We use the original Wikipedia passage as well as passages retrieved by an off-the-shelf system as $\mathbf{C}$. Figure 4 shows the overview of the pipeline.
Seed passage generation.
We randomly sample a Wikipedia article $c$ and use it as our gold reference passage. To produce model-generated seed passages, we diversify the sampled article by paraphrasing and transferring the text to another genre via an LM. Then, we prompt ChatGPT to create a new genre of the text $t$, specifying to include information only found in the Wikipedia article $c$ to maintain factuality. We randomly sample one of the pre-specified genre types (e.g. blog article, question answer pair, tweet, etc.) for each passage. To further diversity in the dataset, we also add examples of question answering (QA) data. In particular, we prompt ChatGPT with two demonstrations from the NaturalQuestions Kwiatkowski et al. (2019) dataset to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles. See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 9.
Error insertion.
In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type). We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types. Therefore, we use ChatGPT and GPT-4 interchangeably for six different types. [^3] We use ChatGPT for the Type 1 and Type 3 while using GPT-4 for other types and insert factual errors one by one into our diversified text. In particular, given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retraining previously inserted errors and original text. The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution.
After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
Post processing.
We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$. At this stage, we also filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types). As a result, we have a training instance $(c,y,y^{*})$, consisting of the gold context (original Wikipedia paragraph) $c$, erroneous LM output $y$, and output with error tags and correct editing $y^{*}$. We also retrieve additional four relevant paragraphs from Wikipedia using Contriever Izacard et al. (2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$. As this process does not involve human annotations, we can easily scale up training instances under an API cost budget.
Statistics of data.
We generated a total of 35,074 training instances, 30,0074 of which are based on Wikipedia passages, and 5,000 are based on QA pairs described above. All of the error types are almost equally distributed (roughly 15% each). We present the error distributions in Appendix Table 10. We conduct a detailed human evaluation in Section 7.2. On average, 3.1 errors are inserted for the Wikipedia passage-based subsets, and 1.4 errors are inserted for the QA pairs.
5.3 Training and Inference
We train a smaller LM on the generated large-scale training data. In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using standard language modeling objective. We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$.
At inference time, we retrieve the top five documents from Wikipedia, [^4] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder. using Contriever-MSMARCO Izacard et al. (2022), and insert them together with an LM output that may include factual errors. The model identifies factual errors, marks phrases or sentences for deletion, and suggests edits for improving the factuality of the text.
6 Experiments
We evaluate Fava and state-of-the-art LMs with and without retrieval on the fine-grained hallucination detection (Section 6.1) and editing tasks (Section 6.2). We further conduct human evaluations on our best models (Section 6.3).
ChatGPT Generations Llama2-Chat 70B Generations
Editor  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
Rt+CGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
Fava (ours) 43.1 25.0 57.1 16.7 70.5 35.3 46.5 78.2 56.2 34.5 27.7 34.5 67.0 43.4 46.5 80.6
Table 2: Overall fine-grained detection performance. CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
6.1 Experiments for Hallucination Detection
Evaluation data.
We use our new benchmark, consisting of 364 annotated passages, [^5] In total we collected 471 annotated passages including our experimental batches, and we excluded earlier batches for evaluations. as our test data. We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.3.
Baselines.
For fine-grained hallucination detection tasks, we test three baselines. ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations. Rt-ChatGPT uses the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt. [^6] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
6.2 Experiments for Editing
Evaluation data and function.
For editing evaluations, we use two open-ended text generation data and evaluate models’ factuality based on two metrics designed to measure the factuality of text. FActScore Min et al. (2023) prompts ChatGPT and GPT 3-davinci 003 to decompose a response into a set of atomic facts and verify factuality for each of them using passages from a designated Wikipedia article. We use the original biography generation task proposed with FActScore for evaluation.
Baselines
For editing experiments, we use ChatGPT and Rt-ChatGPT, as well as Llama2-Chat 13B with and without retrieval (Llama and Rt-Llama, respectively). As in the fine-grained error detection task, we prompt the models with carefully designed prompts and two demonstrations. For all baselines, we retrieve five documents consisting of one document retrieved by string matching based on the entity name and four documents retrieved by Contriever, which are reranked at test time, using a small cross-encoder. [^7] https://huggingface.co/cross-encoder/ms-marco-MiniLM-L-6-v2
6.3 Human Evaluations
Our automatic evaluations may not fully capture the models’ abilities to detect annotated data and editing errors due to the potential subjectivity of annotations and may be affected by the performance of factuality evaluation metrics. We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT. We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents. This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at the individual detection level. Due to the cost of evaluating the factuality of all verification-worthy statements Min et al. (2023), we do not evaluate recall.
7 Results and Analysis
7.1 Results
Detection results.
Table 2 shows the fine-grained detection accuracy of Fava and baselines. We provide the full precision and recall scores in Appendix Tables 12 and 13, respectively. We also provide binary prediction performance.
Table 2 shows that while ChatGPT-based baselines show relatively good binary predictions of existence of factual errors, their fine-grained error type detection performance is significantly lower. Fava shows its strong capabilities of conducting fine-grained error detection, resulting in significant performance improvements from ChatGPT or Ret-ChatGPT on both fine-grained error detection and binary error detection. Specifically, Fava shows high accuracy on error types such as  Contradictory ,  Subjective ,  Entity . On the other hand, its performance on  Invented and  Unverifiable are still limited. Those two error types often require intensive search over many web documents beyond top few passages, while Fava by default only considers top five documents. Our inter-annotator agreements also show that there are more disagreements in those two categories, which might also deflate the performance.
ChatGPT Al-13B Al-7B
No Edit 66.7 42.5 38.8
[…]
Rt-ChatGPT 62.7 -4.0 43.9 +1.4 39.2 +0.4
Llama 52.6 -14.1 22.7 -19.8 18.6 -20.2
Rt-Llama 58.7 -8.0 48.6 +6.1 32.2 -6.6
[…]
Table 3: Overall results on FActScore. Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
Editing results.
Table 3 shows the results on FactScore. Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model size. Retrieval-augmented ChatGPT shows small improvements by editing Alpaca outputs, editing its output with retrieved documents hurts the FActScore. We found that ChatGPT is often confused with different error types and exhibits tendencies of over-predicting contradictory or subjective statements. Among other models, our methods show the largest gains in terms of factuality, showing the effectiveness of editing, in addition to its strong capability of providing fine-grained error types.
Human evaluation results.
Our human evaluation results are shown in Table 4. As in automatic evaluations, our method shows significantly better performance than Rt+ChatGPT on both editing and detection and recognizes more errors than retrieval-augmented ChatGPT. These results further demonstrate the strong capabilities of Fava detecting factual errors in LM outputs.
Model avg. $|\mathbf{E}|$ Detect (%) Edit (%)
Rt+ChatGPT 1.9 23.9 17.2
Fava (ours) 2.4 55.7 63.5
Table 4: Human evaluation results. We show the average number of detected errors and the correctness (%) of the fine-grained types and edits.
7.2 Analysis
Figure 5: Detection scores across different training instance values.
We conduct a set of ablation studies of different aspects of training data creation and inference time design choices, to understand the key factors to build a reliable fine-grained error detection system.
Effects of scale.
We evaluate our model trained on varying numbers of model-generated training data. In particular, we train three Fava variants trained on 10k, 20k, and 30k training instances, and evaluate their performance on fine-grained error detection tasks. Figure 5 shows the results of each variant and demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
Effect of retrieval.
We assess Fava’s editing performance on FActScore on Alpaca 13B outputs through various retrieval methods during inference. Specifically, we examine four distinct retrieval approaches: retrieving the top 1 document; retrieving the top 5 documents; reordering the top 5 documents; retrieving the top 4 documents and mixing them with an introductory paragraph of the target entity (entity matching). We provide the details of the ablation studies in Appendix Section C.4.
Settings FActScore
Alpaca 13B (No Edit) 42.5
Fava (Top 1) 44.2 +1.7
Fava (Top 5) 47.0 +4.5
Fava (Top 5) + reranked 47.7 +5.2
Fava (Top 5) + entity matching 50.1 +7.6
Table 5: FActScore editing results with different retrieval settings.
Table 5 shows the experimental results. Retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%. Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effects of context locations, as reported by prior work Liu et al. (2023a). Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations. Prior work reports that neural retrieval systems can struggle in rare entities Mallen et al. (2022), where term-based retrievers or simple string matching shows more robust performance. FActScore bio consists of extremely low entities, where simple entity matching may help. These results indicate although Fava demonstrates strong capabilities, there’s room for improvements in the retrieval components and how to incorporate many references.
Human evaluation of generated data.
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. We evaluate a generated instance from two aspects: (1) validity—whether the model edits do not violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples. We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
Our analysis revealed that the data generated by our system with iterative insertion obtained an average score of 1.66 for validity assessment and 1.36 for quality assessment. Meanwhile, data generated by one-shot prompting scored an average of 1.1 for validity and 0.9 for quality. Our human evaluation reveals that our LM-generated edits are indeed of higher quality than the one-shot data creation, and provide realistic factual errors.
8 Conclusions and Discussions
This work introduces a new task of automatic hallucination detection, built upon our newly introduced taxonomy, which hierarchically classifies hallucinations in LMs into six categories. We collect the first human-annotated fine-grained hallucination benchmark, consisting of span-level hallucination types, hallucination text, and edits for two LM outputs across multiple domains. As the first step, we introduce a new retrieval-augmented LM trained on synthetically generated fine-grained error editing data following our taxonomies. Empirical results show that Fava significantly outperforms strong baselines by a large margin on both editing and fine-grained detection tasks, while still large room for improvements for automated fine-grained error detection and editing.
Acknowledgement
We thank Sewon Min for fruitful discussions in the early stages. Akari Asai is supported by the IBM Fellowship. We thank Stability AI for providing computing to train and evaluate the LMs in this work, and Microsoft Accelerate Foundation Models Research Program for the access to Azure and OpenAI APIs. We thank our annotators who provide high-quality annotations for this work. This work was funded in part by the DARPA MCS program through NIWC Pacific (N66001-19-2-4031), NSF IIS-2044660, and gifts from AI2.
Appendix
Appendix A Details of Human Annotations
Seed query selections.
Table 6 shows examples of the seed queries for our initial dataset creation.
Dataset Example
WebNLG Explain A.C. Cesena, including information about ground, league.
Instruction-following Explain the differences between New York cheesecakes and Basque cheesecakes in detail.
Open-Assistant Can you tell me how tall the burj khalifa is?
Table 6: Examples of source prompts.
Annotation interface.
Figure 6 shows the interface of our annotations.
Figure 6: Annotation interface.
Details of annotations.
We hired 5 undergrads studying computer science to complete annotations. Each annotator was assigned 60 passages to annotate and had to undergo a 45 minute one on one training session to understand the task and how to navigate the annotation platform. The training session covered an in depth explanation of the six different hallucination types in our taxonomy, included a walk through of how to annotate a passage, and allowed time for annotators to ask any questions.
Examples of annotated results.
Table 7 shows the fine grained hallucination distribution across our annotations. Table 8 shows list of annotated examples from each of our annotation datasets - WebNLG, OpenAssistant, and Instruction Following dataset.
Type Percentage
Entity Errors 28.4%
Relation Errors 2%
Contradictory Statements 6.1%
Invented Statements 3%
Subjective Statements 20.2%
Unverifiable Statements 40.3%
Table 7: Hallucination percentages in our human-annotated data.
Dataset
Annotated Passage Example
WebNLG
A Long Long Way is a novel written by Sebastian Barry. It was first published in 2005 by <entity>Faber and Faber viking press</entity>, a <entity>British American</entity> publishing company. The novel tells the story of Willie Dunne, an Irish soldier who has a strong sense of loyalty to his country, but also feels torn between his Irishness and his loyalty to the British army during World War I. The novel was critically acclaimed and received several literary awards, <contradictory>including the 2005 Dublin Literary Award </contradictory> and was shortlisted for the Man Booker Prize in the same year.
OpenAssistant
<subjective>The most luxurious airline flying within the continental United States is generally regarded as JetBlue Airways</subjective>, particularly for its Mint business class service. Mint offers a private suite experience, featuring fully lie-flat seats, ample legroom and storage, premium food and beverage options, access to onboard WiFi and live TV streaming, and turn-down service complete with duvets and pillows.
Instruction Following
Some of Afghanistan’s secondary cricket home grounds include:
1.<contradictory>Kandahar Cricket Stadium, Kandahar </contradictory>
2.<invented>Mis Ainak Cricket Ground, Logar</invented>
3.<invented>Bost University Cricket Ground, Helmand</invented>
Table 8: Examples from our human annotations for each dataset.
Appendix B More Details of Synthetic Data Creation and Training
Text diversification results.
Table 9 shows the diversified text results.
Seed Passage Diversification Examples
Genre Topic
Output
Article Community Reinvestment Act
The Community Reinvestment Act (CRA) is a federal law that encourages commercial banks and savings associations to help meet the needs of borrowers in low- and moderate-income neighborhoods. The Act was passed in 1977 by Congress to reduce discriminatory credit practices against low-income neighborhoods.
Essay National Dodgeball League
The National Dodgeball League is the only professional dodgeball league in the US, and was founded in 2004. It is headed by Commissioner Edward Prentiss out of Hopkins, MN and is composed of 24 professional teams that are divided into the National and American Dodgeball Conferences. In the past, the tournament has featured teams from Canada, Denmark, UK, Japan, New Zealand and Australia.
Blog Post Red Channels
In the beginning of the 1950s, the "Counterattack" journal published a document, "Red Channels," which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association. Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history. "Red Channels" effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
Social Media Post Sandman
Check out The Sandman - a semi-retired American professional wrestler! He is best known for his career with Extreme Championship Wrestling (ECW), where he held the ECW World Heavyweight Championship a record five times. He also had stints in World Championship Wrestling (WCW), Total Nonstop Action Wrestling (TNA), and World Wrestling Entertainment (WWE). #Wrestling #ECW #WCW #TNA #WWE
News Report I Don’t Mind (James Brown song)
James Brown’s "I Don’t Mind" has made its way up the R&B "Billboard" charts, hitting number four. The track was released in 1961 and also reached number 47 on the Pop "Billboard" charts. Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
Summary Robert Lui
Robert Lui was born on February 23, 1990, and is a professional rugby league footballer from Australia. He plays either as a halfback or five-eighth for the Townsville Blackhawks in the Queensland Cup.
Speech George Sperling
Dear fellow Americans, I am here to propose a solution to improve American Sign Language communication. George Sperling suggests that with a sevenfold reduction in the bandwidth for video transmission, we can achieve this. He even argued that the telephone was originally created for the hearing impaired but it became popularized by the hearing community. Let us not forget our roots and make a change for the better.
Presentation Intro Blissful Ignorance Effect
Have you ever wondered why sometimes people who know less about a product seem to enjoy it more than those who have researched it thoroughly? This phenomenon is called the Blissful Ignorance Effect and it’s a fascinating topic in consumer behavior studies. Our presentation today will explore this effect and why it happens.
Brochure Nidulariaceae Fungi
Explore the fascinating world of Nidulariaceae fungi! This family, found in most ecological regions, includes five different genera, each with its own unique characteristics. With their tiny egg-filled structures, these fungi are a wonder to behold!
Text Message Ashes and Diamonds
Just found out about this book called Ashes and Diamonds by Jerzy Andrzejewski. It’s set during the last few days of WWII. The main character, Maciek, has to kill a Communist soldier. Sounds intense!
Table 9: Text diversification prompts. Instructions for diversification follow the following format: "Given a passage, create a(n) [genre] of 3-6 sentences using only the information present in the passage. Do not include any new information not presented in the passage. Passage: [sampled wikipedia paragraph]"
Error distribution.
Table 10 shows the distribution of error types across all our generated passages.
Percentage
 Entity 21.2%
 Relation 19.9%
 Contradictory 15.3%
 Invented 14.6%
 Subjective 14.1%
 Unverifiable 14.9%
Table 10: Statistics of generated training data
Training Details.
Our base model is Llama2-Chat 7b trained using 4xA40 GPUs. Our training code is based off Open-Instruct Wang et al. (2023a) [^8] https://github.com/allenai/open-instruct/blob/main/scripts/finetune_with_accelerate.sh. Table 11 shows the training hyperparameters.
Precision Epochs Weight Decay Warmup Ratio Learning Rate Max. Seq. Length Batch Size
BFloat16 2 0 0.03 2e-5 2048 128
Table 11: Training hyperparameters.
Appendix C More Results and Analysis
C.1 Detection Results
Overall precision and recall.
Tables 12 and 13 report the precision and recall on our curated datasets.
ChatGPT Generations Llama2-Chat 70B Generations
Model  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 12.1 25.0 50.0 6.7 7.7 0.0 13.0 35.1 17.5 13.2 36.4 15.5 18.8 9.5 19.3 59.2
R+CGPT 19.5 12.5 20.6 3.3 28.6 11.9 17.2 49.7 25.2 16.7 18.8 25.7 14.9 3.2 21.0 60.0
Ours 35.4 20.0 46.2 10.0 75.0 30.0 40.1 69.1 56.8 31.3 23.7 39.6 70.6 46.4 46.1 80.0
Table 12: Fine-grained detection task (Precision). CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
ChatGPT Generations Llama2-Chat 70B
Model  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
CGPT 51.3 33.3 33.3 50.0 7.7 0.0 33.6 87.4 41.8 19.2 21.1 8.5 16.7 20.0 32.4 81.1
R+CGPT 49.9 41.7 33.3 14.3 55.6 22.4 41.7 93.1 50.9 45.8 33.3 19.6 22.2 8.8 41.0 92.5
Ours 55.2 33.3 75.0 50.0 66.7 42.9 55.5 90.0 55.5 38.5 33.3 30.6 63.7 40.6 46.8 81.2
Table 13: Fine-grained detection task (Recall). CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
C.2 Details of Human Evaluation on Model Outputs
We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT. We anonymize the model results to avoid potential biases during evaluations. Annotators count how many of the detection and editing results are correct given extensive web search.
C.3 Manual Analysis on Generated Data for Fava training
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. Prior work introduces entity-centric synthetic data creation Longpre et al. (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed. Therefore, we evaluate not only validity but also quality Table  14 shows examples of generated erroneous data from our synthetic data creation pipeline. We ask human annotators to score each category for each passage either 0, 1, or 2. These scores were defined as follows: a score of 0 indicated that less than half of the inserted errors aligned with the given criteria, a score of 1 signified that at least half of the errors met the established criteria, and a score of 2 indicated that all inserted errors met the criteria. As a comparison, we also generate training data with one-shot generation.
Original Passage
Automatic Data Generations for Training
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive Multiplayer in 1994, the release was ultimately canceled and never reached the market.
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. <invented>The game won the Shooting Star Award in 2020 for its innovative concept.</invented> Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive <entity><delete>Multiplayer</delete><mark>Singleplayer </mark></entity> in 1994, the release was ultimately canceled and never reached the market. <subjective>The game is a masterpiece that revolutionized the arcade industry and will always be remembered as a classic.</subjective>
The knapsack problem is a problem in combinatorial optimization where the goal is to determine the optimal selection of items to maximize the total value within a given weight constraint.
The knapsack problem is a problem in <entity><delete>combinatorial optimization</delete><mark>sequential search</mark></entity> where the goal is to determine the optimal selection of items to maximize the total <entity><delete>value</delete><mark>weight</mark></entity> within a given weight constraint.
Table 14: Examples of generated data from automatic data creation.
C.4 More Ablations
Retrieval ablations.
Settings FActScore
Alpaca 13B (no edits) 42.5
Fava (Top 1) 44.2 +1.7
Fava (Top 5) 47.0 +4.5
Fava (Top 5) + reranked 47.7 +5.2
+ entity search 50.1 +7.6
+ order 51.8 +9.3
Table 15: FActScore editing results with different retrieval settings.
Table 15 shows the full results of the retrieval ablations. We enhance our retrieval process by strategically reordering the retrieval results. This involves rearranging the top 5 passages from our reranked Contriever documents and entity matching document such that the documents with the highest relevance are positioned closest to the text requiring verification. Reordering passages gives 1.7% improvements, further emphasizing the importance of retrieval and careful pipeline design.

### v3

@@
 Abstract
-Large language models (LMs) are prone to generate diverse factually incorrect statements, which are widely called hallucinations.
-Current approaches predominantly focus on coarse-grained automatic hallucination detection or editing, overlooking nuanced error levels.
-In this paper, we propose a novel task—automatic fine-grained hallucination detection—and present a comprehensive taxonomy encompassing six hierarchically defined types of hallucination.
-To facilitate evaluation, we introduce a new benchmark that includes fine-grained human judgments on two LM outputs across various domains.
-Our analysis reveals that ChatGPT and Llama2-Chat exhibit hallucinations in 60% and 75% of their outputs, respectively, and a majority of these hallucinations fall into categories that have been underexplored.
-As an initial step to address this, we train Fava, a retrieval-augmented LM by carefully designing synthetic data generations to detect and correct fine-grained hallucinations.
-On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT on fine-grained hallucination detection by a large margin though a large room for future improvement still exists.
-Fava’s suggested edits also improve the factuality of LM-generated text, resulting in 5-10% FActScore improvements.
+Large language models (LMs) are prone to generate factual errors, which are often called hallucinations.
+In this paper, we introduce a comprehensive taxonomy of hallucinations and argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality.
+We propose a novel task of automatic fine-grained hallucination detection and construct a new evaluation benchmark, FavaBench, that includes about one thousand fine-grained human judgments on three LM outputs across various domains.
+Our analysis reveals that ChatGPT and Llama2-Chat (70B, 7B) exhibit diverse types of hallucinations in the majority of their outputs in information-seeking scenarios.
+We train Fava, a retrieval-augmented LM by carefully creating synthetic data to detect and correct fine-grained hallucinations.
+On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT and GPT-4 on fine-grained hallucination detection, and edits suggested by Fava improve the factuality of LM-generated text.
 [^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
+Fine-grained Hallucination Detection and Editing for Language Models
+[…]
 1 Introduction
@@
 2020) can generate highly fluent and plausible text.
-However, these models are prone to produce hallucinations—factually incorrect generations—which often impedes their deployment in real-world applications Mallen et al.
+However, these models are prone to produce factually incorrect or unverifiable statements, often called hallucinations.
+This impedes their deployment in real-world applications for information-seeking contexts Mallen et al.
 (2022).
-Prior work on hallucinations in natural language generation often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al.
+Prior work on hallucinations in natural language generation (NLG) often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al.
 (2023).
-On the contrary, escalating apprehensions have been articulated with respect to LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al.
-(2023); Huang et al.
+On the contrary, escalating apprehensions have been articulated concerning LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al.
 (2023).
-Figure 1: Overview of fine-grained detection system and the proposed taxonomy for hallucinations.
-Such prevalence of hallucinations in LM outputs has motivated the development of automatic hallucination detection Min et al.
-(2023) or editing systems Gao et al.
+Figure 1: Overview of taxonomy for LM hallucinations and our model, Fava, a fine-grained detection system.
+Figure 2: An overview of our fine-grained hallucination taxonomy. We identify 6 fine-grained types representing diverse hallucinations in LM-generated text.
+Several recent work studies automatic hallucination detection Min et al.
+(2023) or editing outputs Gao et al.
+(2022) to address such LM hallucinations.
+These systems typically categorize hallucinations into simplistic binary distinctions like factual or not factual (Figure 1 top).
+We argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality.
+Entity-level contradictions are usually evident and can be easily rectified with a single reference.
+Conversely, errors involving fabricated entities (e.g., The Messi Diaries in Figure 1) demand thorough verification across multiple sources.
+This underscores the need for a more fine-grained approach to detect hallucination for model development and human verification.
+In this work, we propose automatic fine-grained hallucination detection (Figure 1 bottom), a new task requiring a system to provide precise identification of hallucination sequences, discern different types of hallucinations based on a taxonomy, and suggest refinements.
+We focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters.
+Our taxonomy (Figure 2) hierarchically classifies hallucinations in LM generations into six categories, based on careful pilot studies with NLP experts.
+We construct a new fine-grained hallucination benchmark, FavaBench, by carefully annotating approximately 1,000 responses of three widely used LMs (Llama2-Chat 7B, 70B and ChatGPT [^2] We use gpt-3.5-turbo-0301 throughout this work.) to diverse knowledge-intensive queries.
+Each response is annotated at the span level to identify hallucinations, including erroneous subspace, types, and potential refinements.
+Our analysis reveals all models include at least one hallucination in the majority of their responses (e.g., 70.2% in Llama2 7B and 59.8% in ChatGPT).
+In addition to the widely-studied entity-level errors, other error types like unverifiable sentences make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
+We introduce Fava, a new retrieval-augmented LM that can identify and mark hallucinations at the span level using a unified syntax (Figure 1, left bottom).
+Due to the annotation costs, we design an LM-based synthetic data generation process, and train Fava on the 35k resulting instances.
+We compare Fava with state-of-the-art LMs on fine-grained hallucination detection and editing in LM outputs.
+Fava significantly outperforms ChatGPT with and without external knowledge by 23.7% on fine-grained hallucination detection.
+Furthermore, Fava effectively corrects hallucinations in diverse LM outputs—Fava editing improves FActScore Min et al.
+(2023) of Alapaca 7, 13B, and ChatGPT by 4.4, 9.3 and 3.3%, respectively.
+2 Related Work
+Type Example ChatGPT Llama2-7b Llama2-70b
+ Entity Lionel Andrés Messi was born on June 12 24, 1987. 42.7% 48.6% 55.7%
+ Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 4.7% 3.4% 2.8%
+  Sentence Messi has yet to be captain for the Argentina football team. 18.9% 15.7% 19.9%
+ Invented Messi is known for his famous airplane kick technique. 14.2% 17.4% 9.5%
+ Subjective Lionel Messi is the best soccer player in the world. 9.9% 8.6% 5.8%
+ Unverifiable In his free time, Messi enjoys singing songs for his family. 9.6% 6.3% 6.3%
+Table 1: Distribution of different errors across ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs and examples of hallucinations for each type from our human-annotated data
+Hallucinations in NLG.
+Prior taxonomies proposed for summarization Pagnoni et al.
+(2021) or text simplifications Devaraj et al.
+(2022) often assume the existence of specific source text and focus on faithfulness to the source Ji et al.
+(2023).
+As this is qualitatively different from LMs’ factual hallucinations grounded in world knowledge, we extend prior work and build a new taxonomy for hallucinations in LM outputs.
+Detecting and editing hallucinations in LMs.
+Several recent or concurrent studies propose methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al.
+(2023); Min et al.
+(2023), or focusing on entity-level hallucination editing Gao et al.
 (2022); Chen et al.
 (2023).
-However, those existing systems tend to treat diverse types of factual errors uniformly, often simplifying them into binary categories such as “factual” or “not factual” with a focus on certain domains or specific types of errors.
-While entity-level contradictions (Figure 1 top) are usually evident and can be easily rectified with a single reference, LMs often fabricate entities that lack real-world existence (Figure 1 bottom).
-This variability underscores the need for a more fine-grained approach to hallucination detection for both model development and human verification, and we argue that a detection system should provide precise identification of hallucination sequences, discern different error types, and suggest potential refinements.
-In this work, we introduce a new task, benchmark, and model to facilitate automatic fine-grained hallucination detection (Figure 1).
-We first introduce a new taxonomy of hallucinations and tasks (Section 3), and collect a high-quality benchmark to test models’ ability to provide fine-grained hallucination detection (Section 4).
-We then introduce a new model that is capable of providing span-level hallucination detection and editing (Section 5) as the first step towards building reliable fine-grained hallucination detection systems.
-Our taxonomy (Figure 2; Table 1) hierarchically classifies factual errors in LM generations into six categories.
-We specifically focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters.
-Based on this taxonomy, we propose a new task of fine-grained hallucination detection, in which a system needs to identify locations of factual errors in a response along with fine-grained hallucination types.
-We carefully annotate two widely-used LMs’ responses to diverse instruction-following and knowledge-intensive queries and construct the first fine-grained hallucination annotation benchmark.
-Our benchmark consists of about 400 responses of ChatGPT and Llama2-Chat 70B with span-level annotations of hallucination types based on our proposed taxonomy.
-Our benchmark reveals on average 1.9 and 3.4 hallucinations per response in ChatGPT and Llama2-Chat, respectively.
-It also shows that in addition to the widely-studied entity-level errors, other error types like unverifiable sentences or invented concepts make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
-Due to the expensive costs of annotations, creating training data for this new task is challenging and our preliminary experiments find proprietary LMs such as ChatGPT perform poorly in fine-grained detection.
-To overcome those challenges, we design an LM-based synthetic data generation process, resulting in 35k training instances.
-On this training data, we fine-tune a new retrieval-augmented LM, Fava, which can identify and mark hallucinations at span level using a unified syntax.
-We compare Fava with state-of-the-art LMs including ChatGPT on fine-grained hallucination detection and editing in LM outputs.
-Fava significantly outperforms ChatGPT with and without external knowledge by 22.1% on fine-grained hallucination detection and 14.2% on binary detection.
-Furthermore, Fava effectively fixes factual errors in diverse LM outputs—Fava editing improves FActScore Min et al.
-(2023) of Alapaca 7, 13B, and ChatGPT output by 4.4%, 9.3% and 3.3%, respectively.
-2 Related Work
-Hallucinations in natural language generation.
-Taxonomy of hallucinations in certain downstream tasks such as summarization Pagnoni et al.
-(2021), text simplifications Devaraj et al.
-(2022), or knowledge-grounded dialogues Dziri et al.
-(2022) have been studied.
-Such prior taxonomies often assume the existence of specific source text and focus on faithfulness to the source Ji et al.
-(2023), which is qualitatively different from LMs’ factual hallucinations grounded in world knowledge.
-In addition, prior taxonomies include many task-specific categories that cannot be directly adapted to describe hallucinations in LM generations.
-Detecting and editing hallucinations in LMs.
-Several recent works study methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al.
-(2023); Min et al.
-(2023), or focusing on entity-level factual errors editing Gao et al.
-(2022); Chen et al.
-(2023).
-Recent surveys have attempted to categorize hallucinations in LM generations, but such categorizations often stretch the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023).
-In this work, we propose a fine-grained taxonomy for fact-conflicting hallucinations in long-form text generation to improve factual consistency.
-While prior work often develops factuality verification systems on top of proprietary LMs Chern et al.
-(2023), we introduce a carefully designed training pipeline for a smaller yet more competitive fine-grained detection system.
-Unlike prior qualitative analysis on hallucinations in LM responses on certain knowledge-intensive queries (e.g., open-domain QA, biography generation; Liu et al., 2023b; Min et al., 2023), we annotate responses to diverse queries in multiple domains and reveal hallucination distributions vary across domains.
-Figure 2: An overview of our fine-grained hallucination taxonomy.
-Type Example ChatGPT Llama2
- Entity Lionel Andrés Messi was born on June 12 24, 1987. 46.9% 46.5%
- Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 2.8% 2.5%
- Contradictory Messi has yet to gain captaincy for the Argentina national football team. 12.0% 14.1%
- Invented Messi is also famous for his discovery of the famous airplane kick technique. 4.6% 23.0%
- Subjective Lionel Messi is the best soccer player in the world. 15.0% 8.8%
- Unverifiable In his free time, Messi enjoys singing songs for his family. 18.7% 5.1%
-Table 1: Examples of hallucinations with their percentage in our human-annotated data. We also show the percentage of those errors in ChatGPT and LLama2-Chat-70B outputs.
-Fact verification.
+Recent surveys have attempted to categorize hallucinations in LM generations, often stretching the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023).
+We propose a fine-grained taxonomy for hallucinations in long-form text generation for information-seeking context to improve factual consistency.
+While prior work often develops binary factuality verification systems on top of proprietary LMs Chern et al.
+(2023), we train a fine-grained detection system using our carefully designed data generation pipeline.
+Fact verification for human-written claims.
 While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021).
-Thorne et al.
-(2018) introduce a large scale dataset based on Wikipedia documents for the training and evaluation of fact verification systems.
-Other datasets in information-dense domains like fact-checking against scientific articles (Wadden et al., 2020) and news documents (Wang, 2017) have been proposed.
+Many large-scale datasets have been proposed for fact verification given Wikipedia Thorne et al.
+(2018), scientific articles (Wadden et al., 2020) or news documents (Wang, 2017).
 Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text.
-While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims as well.
+While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims.
 3 Fine-grained Hallucination Detection
-3.1 Focus and Definitions
-In this work, we focus on open-ended text generation given information-seeking queries which often require factual knowledge.
-We define hallucinations as factually incorrect or unverifiable statements given external world knowledge.
+Focus and definitions.
+We focus on open-ended text generation given information-seeking queries which often require factual knowledge.
+We define hallucinations as factual errors or unverified statements given external world knowledge.
 We operationalize world knowledge as documents from the web that are most relevant to the given statement/query according to a search algorithm.
@@
 Based on our definition of hallucinations, we build a hierarchical taxonomy to categorize them into fine-grained error types.
-To develop a hallucination taxonomy for open-ended LM generations, we draw inspiration from error categories in prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022) and introduce new categories to describe more complex errors surfacing in LM generations.
-We conducted a pilot annotation with 9 NLP experts to discuss and refine our taxonomy to ensure good coverage across diverse error types.
+Inspired by prior task-specific taxonomies (Pagnoni et al., 2021; Devaraj et al., 2022), we introduce new categories to describe more complex errors surfacing in LM generations.
+We conducted a pilot annotation with nine NLP experts to ensure good coverage across diverse error types.
 Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination.
@@
 (1a)  Entity : Contradictory entity errors are a sub-category within contradictory statement errors where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
-(1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship described in a statement is incorrect.
-While (1a) often involves errors in nouns, (1b) involves verbs, prepositions, or adjectives, and can be fixed by correcting the incorrect relation.
-(1c)  Contradictory : Contradictory statement errors refer to statements that entirely contradict relevant evidence from the web.
-These are cases where the full sentence is refuted by the information provided in a given reference.
-(2)  Invented : Invented errors refer to statements with concepts that do not exist based on world knowledge.
-These are cases when the LM generates a non-existent or entirely fabricated entity that doesn’t occur in any relevant evidence.
-This does not include fictional characters in creative work such as books or movies.
-(3)  Subjective : Subjective errors refer to an expression or proposition that lacks universal validity and is often influenced by personal beliefs, feelings, opinions, or biases and hence cannot be judged as factually correct.
-Specifically, this represents cases where the LM generates a statement without any factual proposition grounded in relevant evidence.
-(4)  Unverifiable : Unverifiable errors refer to statements that contain factual propositions but cannot be grounded in world evidence.
-These are cases where the LM generates statements with facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal or private matters).
-While  Entity  or  Relation  are often phrase-level and can be fixed by minimal editing erroneous phrases, other error types can be an entire sentence or part of a sentence and should be removed from a response to make it factual.
+(1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship (e.g., verbs, prepositions, or adjectives) in a statement is incorrect.
+(1c)   Sentence : Contradictory sentence errors refer to cases where a full statement entirely contradicts relevant evidence from the web.
+(2)  Invented : Invented errors refer to statements where the LM generates an entirely fabricated entity that doesn’t exist based on world knowledge.
+Fictional entities in creative work aren’t included.
+(3)  Subjective : Subjective errors refer to an expression that lacks universal validity and is often influenced by personal beliefs, opinions, or biases.
+(4)  Unverifiable : This refers to statements where the LM output contains facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal matters).
+[^4] While  Invented and  Unverifiable are highly related,  Invented are the hallucinations where we can verify that some core entities or subjects of the sentences don’t exist.
+While  Entity  or  Relation  are often phrase-level and can be fixed by minimal editing erroneous phrases, other error types can be an entire sentence or part of a sentence and should be removed from a response to make it precisely factual.
 3.2 Tasks and Metrics
-We introduce two new tasks of identifying and editing fine-grained factual errors in LM outputs.
-Given an input query $x$ and a corresponding LM output $y$, our tasks require systems to identify all of the factual errors.
+We introduce the two tasks of identifying and editing fine-grained factual errors in LM outputs.
+Given an input query $x$ and a corresponding LM output $y$, our tasks require systems to identify all of the factual errors in $y$.
 Each error $e$ consists of $(e^{text},e^{type})$, indicating the factually incorrect text spans and their error types among our taxonomies, respectively.
-We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1) and editing factual errors $e^{text}$ accordingly (Task 2).
-Figure 3: Hallucination type distribution of human annotations across 3 datasets and two models.
+We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1; Figure 1 < Entity >) and editing factual errors $e^{text}$ accordingly (Task 2; Figure 1 Barcelona $\rightarrow$ Miami).
+Figure 3: Distribution of hallucination types in ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs across four datasets of diverse information-seeking queries based on our human annotated benchmark.
 Task 1: Fine-grained hallucination detection.
-In the first fine-grained error detection task, the system is expected to identify fine-grained errors in an LM output.
+In this task, the system is expected to identify fine-grained errors in an LM output.
 Due to the subjectivity of span-level annotations, for automatic evaluation, we evaluate systems’ abilities to detect whether an error type $t$ exists in a sentence $s_{i}\in y$.
-Given an output $y$ consisting of $L$ sentences, we assume the availability of gold error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE).
+Given an output $y$ consisting of $L$ sentences, we assume the availability of ground-truth error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE).
+Following the fact verification literature Thorne et al.
+(2018); Schuster et al.
+(2021); Feng et al.
+(2023) of computing precision and recall, we measure those metrics while we extend them for each error type.
 For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
-$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==\texttt{TRUE}]}$ (1)
-$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}==e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}==\texttt{TRUE}]}$ (2)
-Precision indicates how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model.
+$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=\texttt{TRUE}]}$ (1)
+$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}=\texttt{TRUE}]}$ (2)
+Precision indicates the proportion of how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model.
 For final score, we compute the F1 scores averaged over six error types as follows:
 $\frac{1}{|\mathcal{E}|}\sum_{t\in\mathcal{E}}\frac{2\times\text{Prec}^{t}\times\text{Recall}^{t}}{\text{Prec}^{t}+\text{Recall}^{t}}$ (3)
-Fine-grained error detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not.
-This is similar to FActScore Min et al.
-(2023), where we predict and evaluate factuality at the sentence level without extracting atomic facts.
+Fine-grained detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not.
+[^5] This is similar to FActScore Min et al.
+(2023), while we predict and evaluate factuality at the sentence level without extracting atomic facts.
 Task 2: Hallucination editing.
-As shown in Table 1, some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable.
+Some hallucinations can be fixed with minimal span-level edits, while others need to be removed or marked as unverifiable.
 In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output.
-We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore or QAFactEval Fabbri et al.
-(2022).
-as follows:$f(\hat{y})-f(y)$, where $f$ indicates an evaluation function for factuality while $f(\hat{y})$ indicates edited output.
-4 Benchmark
-4.1 Human Annotation on Hallucinations
-To facilitate model development in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are, we carefully annotate the first fine-grained hallucination detection benchmark.
-Our benchmark consists of multi-way fine-grained annotations on two LM responses to queries in multiple domains.
+We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore Min et al.
+(2023), as: $f(\hat{y})-f(y)$, where $f$ indicates an estimated factuality scores and $\hat{y}$ indicates edited output.
+4 Benchmark: FavaBench
+We create FavaBench to facilitate model development and evaluation in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are.
+FavaBench consists of multi-way fine-grained annotations on three LM responses to queries in multiple domains.
 Source prompts.
-To evaluate LMs’ factuality in the responses to diverse queries, we collect 150 information-seeking queries, spanning three different categories.
-First, we sample 50 knowledge-intensive queries from the Open Assistant Köpf et al.
-(2023) dataset, by prompting GPT-4 to judge whether each query from the dataset requires world knowledge or not.
-We, the authors further compose 50 instruction-following queries that require more complex reasoning and knowledge.
-Lastly, to enable more controlled evaluations and analysis, we synthetically create 50 queries that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al.
-(2017) using a template.
-See examples in each category in Appendix Table 6.
-Annotation details.
-We prompt ChatGPT Ouyang et al.
-(2022) and Llama2-Chat 70B Touvron et al.
-(2023) in a zero-shot manner and collect 300 responses to our diverse information-seeking queries.
-We recruited 15 students (five undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy.
-Due to the complexity of the task, we conducted 45-minute in-person and virtual training sessions with them.
-Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation.
-For each LM response, we assign two annotators.
-We present model-generated responses, input information-seeking queries, and the top 10 web search documents returned by Google Search given the original input queries.
-We ask human annotators to first mark a span, annotate a category of error, and suggest an edit to fix the errors for the  Entity  and  Relation  types.
-Our annotation interface and more details of annotations are available in Appendix A.
+The source prompts include a collection of 200 information-seeking queries, spanning four different data sources. See examples in Appendix Table 4.
+  • Knowledge-intensive queries sampled from the Open Assistant dataset Köpf et al. (2023), by prompting GPT-4 to judge whether each query from the dataset requires world knowledge.
+  • A sample of Open QA prompts from the No Robots dataset Rajani et al. (2023).
+  • Instruction-following prompts requiring more reasoning and knowledge curated by the authors.
+  • Synthetically-created prompts 50 prompts that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template.
+Annotation details and quality.
+We obtain responses to the collected source prompts using ChatGPT (gpt-3.5-turbo-0301;  Ouyang et al.
+2022) Llama2-Chat 7B and Llama2-Chat 70B Touvron et al.
+(2023) in a zero-shot manner and collect 600 responses to our diverse information-seeking prompts.
+We recruited 20 students (ten undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy.
+Each instance is annotated by two annotators who completed 45-minute in-person and virtual training sessions.
+[^6] Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation.
+Our annotation interface and details are in Appendix A.
+To validate our annotation quality, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in detection across passages, with 75.1% agreement in detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages.
+Appendix Table 6 shows examples of our annotation results.
 Analysis on annotated data.
 Figure 3 presents a detailed breakdown of distributions across fine-grained categories in the three domains.
-60.8% and 74.7% of the responses of ChatGPT and Llama2-Chat include at least one factual hallucination, respectively.
-Entity  is the most widely recognized error type, making up 46.3% of the detected errors.
-Yet, there are diverse types of errors prevalent like invented statements or contradictory statements which make up 16.8% and 15.4% of the errors detected respectively.
-We found that the error distributions vary across different source prompts and LMs.
-Invented  are more common in Llama2-Chat 70B.
-We observe fewer errors in Open Assistant subsets, possibly because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How tall is the Burj Khalifa?), which is often memorized by powerful LMs Mallen et al.
+59.8%, 70.2% and 64.9% of the responses of ChatGPT, Llama2-Chat 7B and Llama2-Chat 70B include at least one hallucination, respectively.
+Entity  is the most widely recognized error type, making up 48.3% of the detected errors.
+Yet, there are diverse types of errors prevalent like invented statements or contradictory statements, which make up 14.1% and 18.1% of the detected errors, respectively.
+The error distributions vary across different source prompts and LMs.
+Invented  are more common in Llama2-Chat models.
+Fewer errors in No Robots subsets than other subsets may be because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How long was the Revolutionary War?), which is often memorized by LMs Mallen et al.
 (2022).
-On the other hand, the newly annotated instructions and WebNLG data show a higher percentage of errors.
-Figure 4: Overview of data creation process in Fava.
-Our annotators also agreed that our taxonomy covers most of the error types in LM responses.
-Furthermore, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in error detection across passages, with 75.1% agreement in error detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages.
-Appendix Table 8 shows examples of our annotations.
+Figure 4: Overview of high-quality synthetic data generation process in Fava.
+Fava leverages powerful instruction-tuned models to carefully insert errors into factually accurate statements and produces diverse error types based on our proposed taxonomy.
 5 Model: Fava
-Our benchmark analysis reveals the urgent need to study and detect such diverse types of hallucinations.
-We introduce a new retrieval-augmented LM, Fava (FAct Vericaton with Augmentation).
-Fava is trained on high-quality synthetic training data, and at inference, it identifies and fixes fine-grained hallucinations, incorporating retrieved knowledge.
-5.1 Overview
+We introduce Fava (FAct Vericaton with Augmentation), a model for fine-grained hallucinations detections and editing.
+Fava is trained on high-quality synthetic training data to identify hallucinations, incorporating retrieved knowledge.
+Overview.
 Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$.
 $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$.
-Subsequently, the editing model detects and if possible edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$.
+Subsequently, the editing model detects and, if possible, edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$.
 $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1.
 While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2).
-Reliance on black-box proprietary API models also hurt reproducible.
-Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.2) and fine-tune a powerful expert LM consisting of 7 billion parameters (Section 5.3).
-5.2 Training Data Creation
-To train our $\mathcal{M}_{edit}$ model, we require a large number of erroneous LM outputs $y$ paired with edited versions $y^{*}$ with fine-grained error types and edits.
+Reliance on such API models also hurt reproducibility.
+Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.1) and fine-tune a smaller yet powerful expert LM on the generated data (Section 5.2).
+5.1 Synthetic Training Data Curation
+To train our $\mathcal{M}_{edit}$ model, we require a large number of training instances, each of which includes $(c,y,y^{*})$, where $c$ is the gold context (original Wikipedia paragraph), $y$ is an erroneous LM output, and $y^{*}$ is an output with error tags and correct editing based on $c$.
+Fava leans to take $(c,y)$ and generate $y^{*}$.
+Yet, such data is not publicly available, and our annotation is expensive to scale up.
 Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al.
@@
 (2023b); Asai et al.
-(2023), we introduce a new data creation method grounded on our fine-grained hallucination taxonomies.
-We prompt GPT-4 Achiam et al.
-(2023) and ChatGPT to diversify and noise a Wikipedia passage by inserting different types of hallucinations.
-In particular, our data creation pipeline consists of three steps: seed passage generations, error insertions, and postprocessing.
-The first two processes create an input sequence with perturbed diverse error annotations, and postprocessing remaps those special tokens to generate the erroneous output $y$ and output with fine-grained detection results $y^{*}$.
-We use the original Wikipedia passage as well as passages retrieved by an off-the-shelf system as $\mathbf{C}$.
+(2023), we introduce a new data creation method tailored to our fine-grained hallucination taxonomies.
+Specifically, our data creation pipeline consists of three steps: Step 1: seed passage generations that generates a large, diverse collection of seed passages $c$, Step 2: error insertions which inserts factually incorrect statements informed by our our hallucination taxonomy to generate erronous passages $y$, and Step 3: post-processing which prunes and curate the final training examples.
 Figure 4 shows the overview of the pipeline.
 Seed passage generation.
-We randomly sample a Wikipedia article $c$ and use it as our gold reference passage.
-To produce model-generated seed passages, we diversify the sampled article by paraphrasing and transferring the text to another genre via an LM.
-Then, we prompt ChatGPT to create a new genre of the text $t$, specifying to include information only found in the Wikipedia article $c$ to maintain factuality.
-We randomly sample one of the pre-specified genre types (e.g.
-blog article, question answer pair, tweet, etc.) for each passage.
-To further diversity in the dataset, we also add examples of question answering (QA) data.
-In particular, we prompt ChatGPT with two demonstrations from the NaturalQuestions Kwiatkowski et al.
-(2019) dataset to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles.
+The goal in this step is to collect and generate a large, diverse range of seed passages to ensure Fava can handle a wide range of text types while maintaining control over factuality.
+Later on, errors are inserted into these seed passages to generate erroneous text, $y$.
+We base our seed passage collection on Wikipedia articles and available Question Answering (QA) datasets Kwiatkowski et al.
+(2019) since they are both good sources to evaluate factuality.
+However, the text genre is quite limited.
+To diversify the passage genres, we paraphrase Wikipedia passages into different genres to enable Fava to be adaptable and effective across various textual formats.
+More specifically, we randomly sample 35,074 Wikipedia articles $c$ and use them as our gold reference passages.
+We then randomly sample one of the pre-specified genre types (e.g.
+blog article, tweet, etc.) for each passage and use ChatGPT to paraphrase the original passage into the given genre.
+In addition, we use 5,000 examples from QA datasets to further diversify the passage collections.
+In particular, we prompt ChatGPT with two QA demonstrations to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles.
 See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 7.
 Error insertion.
+The goal of this step is to simulate erroneous LM generation $y$, which are passages with factual errors.
+To control error distributions in the training data and incorporate our hallucination taxonomy, we prompt LMs to insert errors from our taxonomy one by one.
 In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type).
@@
 Therefore, we use ChatGPT and GPT-4 interchangeably for six different types.
-[^3] We use ChatGPT for the Type 1 and Type 3 while using GPT-4 for other types and insert factual errors one by one into our diversified text.
-In particular, given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retraining previously inserted errors and original text.
+[^7] We use ChatGPT (gpt-3.5-turbo-0301) for the Type 1 and Type 3 while using GPT-4 for other types.
+Given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retaining previously inserted errors and original text.
 The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution.
 After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
-Post processing.
+Post-processing.
 We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$.
-At this stage, we also filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types).
-As a result, we have a training instance $(c,y,y^{*})$, consisting of the gold context (original Wikipedia paragraph) $c$, erroneous LM output $y$, and output with error tags and correct editing $y^{*}$.
-We also retrieve additional four relevant paragraphs from Wikipedia using Contriever Izacard et al.
+We filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types).
+We also retrieve additional four relevant Wikipedia paragraphs using Contriever-MSMARCO Izacard et al.
 (2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$.
-As this process does not involve human annotations, we can easily scale up training instances under an API cost budget.
 Statistics of data.
-We generated a total of 35,074 training instances, 30,0074 of which are based on Wikipedia passages, and 5,000 are based on QA pairs described above.
-All of the error types are almost equally distributed (roughly 15% each).
-We present the error distributions in Appendix Table 10.
-We conduct a detailed human evaluation in Section 7.2.
-On average, 3.1 errors are inserted for the Wikipedia passage-based subsets, and 1.4 errors are inserted for the QA pairs.
+After analyzing ablations with varying training instance sizes (Section 7.2), we generated 35,074 training instances,  [^8] While we see performance improvements up to 30k, the API costs for training data generation were already \$3k, reaching our initial budget.
+30,0074 based on Wikipedia passages, and 5,000 based on QA pairs.
+All of the error types are almost equally distributed ($\sim$15 % each) with 3.1 errors inserted for Wikipedia subsets, and 1.4 errors inserted for the QA pairs on average.
+We present the error distributions in Appendix Table 8 and a detailed human evaluation in Appendix C.4.
 5.2 Training and Inference
 We train a smaller LM on the generated large-scale training data.
-In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using standard language modeling objective.
+In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using the standard language modeling objective.
 We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$.
+We empirically found initializing Fava with this model gives slightly better performance than the pre-trained Llama2.
 At inference time, we retrieve the top five documents from Wikipedia, [^9] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder.
-using Contriever-MSMARCO Izacard et al.
-(2022), and insert them together with an LM output that may include factual errors.
-The model identifies factual errors, marks phrases or sentences for deletion, and suggests edits for improving the factuality of the text.
+using Contriever-MSMARCO, and input them together with an LM output that may include hallucinations.
+The model identifies hallucinations, marks phrases or sentences for deletion, and suggests edits.
 6 Experiments
-We evaluate Fava and state-of-the-art LMs with and without retrieval on the fine-grained hallucination detection (Section 6.1) and editing tasks (Section 6.2).
-We further conduct human evaluations on our best models (Section 6.3).
-ChatGPT Generations Llama2-Chat 70B Generations
-Editor  ent  rel  con  inv  subj  unv OA Bi.  ent  rel  con  inv  subj  unv OA Bi.
-CGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
-Rt+CGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
-Fava (ours) 43.1 25.0 57.1 16.7 70.5 35.3 46.5 78.2 56.2 34.5 27.7 34.5 67.0 43.4 46.5 80.6
-Table 2: Overall fine-grained detection performance. CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
+Generator: ChatGPT Generator: Llama2-Chat 70B
+Editor  ent  rel  con  inv  subj  unv OA Bi  ent  rel  con  inv  subj  unv OA Bi
+ChatGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
+Rt+ChatGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
+GPT4 38.6 16.6 17.9 22.2 50.0 17.2 34.2 60.8 55.5 60.0 21.2 15.4 2.0 25.0 42.5 74.2
+Fava (ours) 54.5 25.0 66.7 16.7 70.5 35.3 48.1 79.6 57.3 34.5 27.7 52.2 31.25 43.4 47.2 80.3
+Table 2: Fine-grained detection F1. OA and Bi indicates overall and binary predictions.
 6.1 Experiments for Hallucination Detection
 Evaluation data.
-We use our new benchmark, consisting of 364 annotated passages, [^5] In total we collected 471 annotated passages including our experimental batches, and we excluded earlier batches for evaluations.
+We use our new benchmark, consisting of 902 annotated passages, [^10] We collected 1010 annotated passages including our experimental batches, excluding earlier batches for evaluations.
 as our test data.
@@
 ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations.
-Rt-ChatGPT uses the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt.
+GPT4 (gpt-4) does the exact same but prompts GPT4 instead.We were unable to include GPT4 with retrieval in our baseline due to high costs.
+Since Fava uses retrieval augmentation, for comparability we include Rt+ChatGPT.
+This baseline prompts ChatGPT using the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt.
 [^11] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
-6.2 Experiments for Editing
-Evaluation data and function.
-For editing evaluations, we use two open-ended text generation data and evaluate models’ factuality based on two metrics designed to measure the factuality of text.
-FActScore Min et al.
-(2023) prompts ChatGPT and GPT 3-davinci 003 to decompose a response into a set of atomic facts and verify factuality for each of them using passages from a designated Wikipedia article.
-We use the original biography generation task proposed with FActScore for evaluation.
-Baselines
-For editing experiments, we use ChatGPT and Rt-ChatGPT, as well as Llama2-Chat 13B with and without retrieval (Llama and Rt-Llama, respectively).
-As in the fine-grained error detection task, we prompt the models with carefully designed prompts and two demonstrations.
+6.2 Experiments for Hallucination Editing
+Evaluation data.
+For editing evaluations, we use open-ended text generation data and evaluate models’ factuality based on a metric designed to measure the factuality of text, FActScore Min et al.
+(2023).
+This model-based metric prompts ChatGPT and GPT 3 (davinci-003) to decompose a response into a set of atomic facts and verify factuality for each using passages from a designated Wikipedia article.
+We use the original biography generation task proposed with FActScore.
+Baselines.
+For editing experiments, we use ChatGPT and Rt-ChatGPT, as well as Llama2-Chat 13B with and without retrieval (Llama and Rt-Llama, respectively), and prompt them with carefully designed prompts and two demonstrations.
 For all baselines, we retrieve five documents consisting of one document retrieved by string matching based on the entity name and four documents retrieved by Contriever, which are reranked at test time, using a small cross-encoder.
@@
 6.3 Human Evaluations
-Our automatic evaluations may not fully capture the models’ abilities to detect annotated data and editing errors due to the potential subjectivity of annotations and may be affected by the performance of factuality evaluation metrics.
+Our automatic evaluations may not fully capture the models’ abilities to detect and refine hallucinations, due to the potential subjectivity of our annotations or the performance of the factuality evaluation metric.
 We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT.
@@
 Detection results.
-Table 2 shows the fine-grained detection accuracy of Fava and baselines.
-We provide the full precision and recall scores in Appendix Tables 12 and 13, respectively.
-We also provide binary prediction performance.
-Table 2 shows that while ChatGPT-based baselines show relatively good binary predictions of existence of factual errors, their fine-grained error type detection performance is significantly lower.
-Fava shows its strong capabilities of conducting fine-grained error detection, resulting in significant performance improvements from ChatGPT or Ret-ChatGPT on both fine-grained error detection and binary error detection.
-Specifically, Fava shows high accuracy on error types such as  Contradictory ,  Subjective ,  Entity .
-On the other hand, its performance on  Invented and  Unverifiable are still limited.
-Those two error types often require intensive search over many web documents beyond top few passages, while Fava by default only considers top five documents.
-Our inter-annotator agreements also show that there are more disagreements in those two categories, which might also deflate the performance.
-ChatGPT Al-13B Al-7B
+Table 2 shows the fine-grained detection accuracy, namely F1 scores, and binary prediction F1 of Fava and baselines.
+We provide the full precision and recall scores in Appendix Tables 11 and 12, respectively.
+Fava significantly outperforms ChatGPT, Ret-ChatGPT, or GPT-4 on both fine-grained error detection and binary error detection.
+Fava shows high accuracy on error types such as   Sentence ,  Subjective ,  Entity .
+On the other hand, its performance on  Invented and  Unverifiable is still limited.
+Those two error types often require intensive search over many documents beyond the top few ones, while Fava by default only considers the top five documents.
+Generator
+Editor ChatGPT Al-13B Al-7B
 No Edit 66.7 42.5 38.8
@@
 […]
-Table 3: Overall results on FActScore. Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
+Table 3: Overall Editing results on FActScore Min et al. (2023). Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
 Editing results.
-Table 3 shows the results on FactScore.
-Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model size.
-Retrieval-augmented ChatGPT shows small improvements by editing Alpaca outputs, editing its output with retrieved documents hurts the FActScore.
-We found that ChatGPT is often confused with different error types and exhibits tendencies of over-predicting contradictory or subjective statements.
-Among other models, our methods show the largest gains in terms of factuality, showing the effectiveness of editing, in addition to its strong capability of providing fine-grained error types.
+Table 3 shows the results of the editing task on FActScore.
+Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model in size.
+We found that ChatGPT is often confused with different errors and has a tendency to over-predict contradictory or subjective statements.
+Fava shows the largest gains in FActScore, showing the effectiveness of editing, in addition to error type detection.
 Human evaluation results.
-Our human evaluation results are shown in Table 4.
-As in automatic evaluations, our method shows significantly better performance than Rt+ChatGPT on both editing and detection and recognizes more errors than retrieval-augmented ChatGPT.
+Our human evaluation results are shown in Appendix Table 14.
+Fava shows significantly better performance than Rt+ChatGPT on both editing and detection scoring 46.3% and 31.8% more than Rt+ChatGPT on each task respectively.
+Furthermore, Fava recognizes more errors, 2.4 errors per passage, than retrieval-augmented ChatGPT, 1.9 errors per passage.
 These results further demonstrate the strong capabilities of Fava detecting factual errors in LM outputs.
-Model avg. $|\mathbf{E}|$ Detect (%) Edit (%)
-Rt+ChatGPT 1.9 23.9 17.2
-Fava (ours) 2.4 55.7 63.5
-Table 4: Human evaluation results. We show the average number of detected errors and the correctness (%) of the fine-grained types and edits.
 7.2 Analysis
-Figure 5: Detection scores across different training instance values.
-We conduct a set of ablation studies of different aspects of training data creation and inference time design choices, to understand the key factors to build a reliable fine-grained error detection system.
-Effects of scale.
-We evaluate our model trained on varying numbers of model-generated training data.
-In particular, we train three Fava variants trained on 10k, 20k, and 30k training instances, and evaluate their performance on fine-grained error detection tasks.
-Figure 5 shows the results of each variant and demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
+Effects of data scale.
+We assess our model on varying sizes of synthetic training data (10k, 20k, and 30k instances) and examine their performance on fine-grained hallucination detection.
+Figure 4(b) demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
 Effect of retrieval.
-We assess Fava’s editing performance on FActScore on Alpaca 13B outputs through various retrieval methods during inference.
-Specifically, we examine four distinct retrieval approaches: retrieving the top 1 document; retrieving the top 5 documents; reordering the top 5 documents; retrieving the top 4 documents and mixing them with an introductory paragraph of the target entity (entity matching).
-We provide the details of the ablation studies in Appendix Section C.4.
-Settings FActScore
-Alpaca 13B (No Edit) 42.5
-Fava (Top 1) 44.2 +1.7
-Fava (Top 5) 47.0 +4.5
-Fava (Top 5) + reranked 47.7 +5.2
-Fava (Top 5) + entity matching 50.1 +7.6
-Table 5: FActScore editing results with different retrieval settings.
-Table 5 shows the experimental results.
-Retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%.
-Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effects of context locations, as reported by prior work Liu et al.
-(2023a).
+We assess Fava’s editing performance on Alpaca 13B outputs with varying retrieved evidence: top 1 document; top 5 documents; reordering the top 5 documents; top 4 documents and mix them with an introductory paragraph of the target entity (entity matching).
+Table 4(a) shows the experimental results, and more detailed results are available in Appendix Table  16.
+We found that retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%.
+Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effect of context ordering, as reported by prior work Liu et al.
+(2023).
 Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations.
-Prior work reports that neural retrieval systems can struggle in rare entities Mallen et al.
-(2022), where term-based retrievers or simple string matching shows more robust performance.
-FActScore bio consists of extremely low entities, where simple entity matching may help.
-These results indicate although Fava demonstrates strong capabilities, there’s room for improvements in the retrieval components and how to incorporate many references.
-Human evaluation of generated data.
-We conduct human evaluations on 50 generated data to assess the automatic data creation quality.
-We evaluate a generated instance from two aspects: (1) validity—whether the model edits do not violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples.
-We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
-Our analysis revealed that the data generated by our system with iterative insertion obtained an average score of 1.66 for validity assessment and 1.36 for quality assessment.
-Meanwhile, data generated by one-shot prompting scored an average of 1.1 for validity and 0.9 for quality.
-Our human evaluation reveals that our LM-generated edits are indeed of higher quality than the one-shot data creation, and provide realistic factual errors.
-8 Conclusions and Discussions
-This work introduces a new task of automatic hallucination detection, built upon our newly introduced taxonomy, which hierarchically classifies hallucinations in LMs into six categories.
-We collect the first human-annotated fine-grained hallucination benchmark, consisting of span-level hallucination types, hallucination text, and edits for two LM outputs across multiple domains.
-As the first step, we introduce a new retrieval-augmented LM trained on synthetically generated fine-grained error editing data following our taxonomies.
+These results indicate although Fava demonstrates strong capabilities, there’s room for improvement in retrieving and incorporating many references.
+Settings Score
+No Edit 42.5
+(Top 1) 44.2 +1.7
+(Top 5) 47.0 +4.5
+Reranked 47.7 +5.2
+Entity 50.1 +7.6
+(a) Retrieval
+(b) Training Size
+Figure 5: (a): FActScore results with different retrieval settings. (b): Detection scores across different training instance values.
+8 Conclusions
+We introduce a new task of automatic hallucination detection, built upon our newly introduced taxonomy, hierarchically classifying hallucinations in LMs into six categories.
+We collect the first human-annotated fine-grained hallucination benchmark, and introduce a new retrieval-augmented LM for fine-grained error editing data following our taxonomies.
 Empirical results show that Fava significantly outperforms strong baselines by a large margin on both editing and fine-grained detection tasks, while still large room for improvements for automated fine-grained error detection and editing.
+Limitations
+This work focuses on hallucinations in LM generations in an information-seeking context, where factual correct matters, and we build a new taxonomy, benchmark, and model for this.
+Errors in other contexts (e.g., commonsense reasoning, numerical reasoning) are beyond our focus.
+Our training data uses Wikipedia passages as gold context, and Fava uses Wikipedia passages as references at inference time.
+Yet, Wikipedia may not cover all of the necessary information to verify the factuality of LM-generated statements Asai and Choi (2021), and it is an open question how well Fava performs when it uses other types of documents as reference documents.
+Our analysis and findings have been based only on the English language and our scope is limited to information-seeking scenarios.
+Our model, Fava, and benchmark, FavaBench, were also trained on English text data.
+To understand fine-grained hallucinations in diverse languages or scenarios, more analysis and experiments need to be done.
+Our annotations were collected with undergraduate and graduate computer science students who attended 45-minute training sessions to guarantee quality.
+Furthermore, annotations were reviewed for quality by us authors and annotator agreement was calculated to ensure understanding within annotators.
+Yet, given the complexity of our annotation tasks, there may be subjectivity or errors in annotations.
+We will release our data publicly to enable further investigations.
+It is difficult to guarantee that our proposed hallucination taxonomy is all-encompassing.
+To address this concern, we asked annotators to include hallucination types not found in our taxonomy during the annotation process.
+We received no additions to our taxonomy during this, indicating our taxonomy is thorough.
+We evaluate Fava to determine feasibility and quality. Although this doesn’t assure quality outputs from Fava consistently, the positive results indicate quality within our defined scope and setup.
+Ethics Statement and Broader Impacts
+LM-generated text is prone to hallucinations, which can result in severe real-world risks such as spreading misinformation.
+In this work, we attempt to understand such prevalent hallucinations by providing a new taxonomy, a human-annotated benchmark, and a new expert LM trained for fine-grained hallucinations.
+We hope our work will inspire future work and encourage more work focusing on fine-grained hallucination detections.
+Yet, our resulting models are not perfect and may overlook hallucinations in LM outputs and we acknowledge such limitations in our paper.
+For evaluating the detection capabilities of our model, we use our annotation data.
+Our annotation collection process did not involve gathering private information or data from our annotators.
+Additionally, of the four datasets we used for annotation prompts, three are publicly available datasets containing no personal information.
+The remaining prompt dataset was curated by us authors and does not include personal information.
+Our annotators were all paid for their time in completing annotations at the rate of \$20 per hour for 10 hours of work across a week.
+They all disclosed the payment amount and estimated time before beginning annotations.
+None of their data was collected while constructing the annotation dataset.
+For evaluating the editing capabilities of our model, we use FActScore Min et al.
+(2023) which has publicly available code and data and is based on Wikipedia.
 Acknowledgement
@@
 Open-Assistant Can you tell me how tall the burj khalifa is?
+No Robots When was Samsung founded?
 Table 4: Examples of source prompts.
@@
 Details of annotations.
-We hired 5 undergrads studying computer science to complete annotations.
+We hired 10 undergrads studying computer science to complete annotations.
 Each annotator was assigned 60 passages to annotate and had to undergo a 45 minute one on one training session to understand the task and how to navigate the annotation platform.
 The training session covered an in depth explanation of the six different hallucination types in our taxonomy, included a walk through of how to annotate a passage, and allowed time for annotators to ask any questions.
+Annotators were given an instruction document outlining the task, details on the hallucination type, details on how to navigate the annotation platform, and payment details which they looked over during their training session before starting annotation work.
+Figure 7 shows the instructions on using the annotation platform and payment details provided to annotators.
+Figure 7: Annotation instructions and payment details provided to annotators.
 Examples of annotated results.
@@
 Blog Post Red Channels
-In the beginning of the 1950s, the "Counterattack" journal published a document, "Red Channels," which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association.
+In the beginning of the 1950s, the ”Counterattack” journal published a document, ”Red Channels,” which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association.
 Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history.
-"Red Channels" effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
+”Red Channels” effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
 Social Media Post Sandman
@@
 News Report I Don’t Mind (James Brown song)
-James Brown’s "I Don’t Mind" has made its way up the R&B "Billboard" charts, hitting number four.
-The track was released in 1961 and also reached number 47 on the Pop "Billboard" charts.
+James Brown’s ”I Don’t Mind” has made its way up the R&B ”Billboard” charts, hitting number four.
+The track was released in 1961 and also reached number 47 on the Pop ”Billboard” charts.
 Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
@@
 Table 7: Text diversification prompts.
-Instructions for diversification follow the following format: "Given a passage, create a(n) [genre] of 3-6 sentences using only the information present in the passage.
+Instructions for diversification follow the following format: ”Given a passage, create a(n) [genre] of 3-6 sentences using only the information present in the passage.
 Do not include any new information not presented in the passage.
-Passage: [sampled wikipedia paragraph]"
+Passage: [sampled wikipedia paragraph]”
 Error distribution.
@@
  Relation 19.9%
- Contradictory 15.3%
+  Sentence 15.3%
  Invented 14.6%
@@
 C.1 Detection Results
+Llama2-Chat 7B F1 results
+Table  10 reports the detection results for the Llama2-Chat 7B generations on our curated datasets.
+Generator: Llama2-Chat 7B
+Editor  ent  rel  con  inv  subj  unv OA Bi
+ChatGPT 25.2 12.6 16.1 15.0 12.4 12.8 26.5 63.4
+Rt+ChatGPT 35.5 17.5 13.2 22.2 10.9 14.6 32.9 70.4
+GPT4 45.4 23.1 28.8 15.4 3.4 28.6 47.3 72.1
+Fava (ours) 58.3 38.7 24.2 58.9 31.25 44.4 39.6 79.9
+Table 10: Fine-grained detection F1. OA and Bi indicates overall and binary predictions.
 Overall precision and recall.
-Tables 12 and 13 report the precision and recall on our curated datasets.
+Tables 11 and 12 report the precision and recall on our curated datasets.
 ChatGPT Generations Llama2-Chat 70B Generations
@@
 Table 12: Fine-grained detection task (Recall). CGPT indicates ChatGPT. “OA” indicates overall accuracy and “Bi.” indicates binary predictions accuracy.
-C.2 Details of Human Evaluation on Model Outputs
+C.2 Benchmark Prompts
+Figure C.2 shows the prompt used for baseline models when evaluating on the detection benchmark.
+Given a passage with factual errors, identify any <entity>, <relation>,
+<contradictory>, <subjective>, <unverifiable> or <invented> errors in the passage and
+add edits for <entity> and <relation> errors by inserting additional <mark></mark> or
+<delete></delete> tags  to mark and delete. If there are no errors, return the passage
+with no tags. Any changes to the original passage should be marked in <> tags. Below
+are the error definitions followed by examples of what you need to follow.
+Definitions:
+1. entity errors (<entity>): a small part of a sentence, often an entity (e.g.,
+location name), is incorrect (usually 1-3 words). Entity errors often involve
+noun phrases or nouns.
+2. relational error (<relation>): a sentence is partially incorrect as a small part
+(usually 1 - 3 words). Relational errors often involve verbs and are often the
+opposite of what it should be.
+3. contradictory sentence error (<contradictory>): a sentence where the
+entire sentence is contradicted by the given reference, meaning the sentence
+can be proven false due to a contradiction with information in the passage.
+4. invented info error (< invented >): these errors refer to entities that are
+not known or do not exist. This does not include fictional characters in books or movies.
+invented errors include phrases or sentences which have unknown entities or
+misleading information.
+5. subjective sentence (<subjective>): an entire sentence or phrase that is subjective
+and cannot be verified, so it should not be included.
+6. unverifiable sentence (<unverifiable>): a sentence where the whole sentence or
+phrase is unlikely to be factually grounded although it can be true, and the sentence
+cannot be confirmed nor denied using the reference given or internet search, it is
+often something personal or private and hence cannot be confirmed.
+Follow the given example exactly, your task is to create the edited completion
+with error tags <>:
+Passage: Marooned on Mars is a science fiction novel aimed at a younger audience.
+It was written by Andy Weir and published by John C. Winston Co. in 1952, featuring
+illustrations by Alex Schomburg. It ended up having a readership of older boys despite
+efforts for it to be aimed at younger kids. The novel inspired the famous Broadway
+musical "Stranded Stars," which won six Tony Awards. The novel tells a story of being
+stranded on the Purple Planet. I wish the novel had more exciting and thrilling plot
+twists.
+Reference: Marooned on Mars is a juvenile science fiction novel written by American
+writer Lester del Rey. It was published by John C. Winston Co. in 1952 with illustrations
+by Alex Schomburg.
+Edited: Marooned on Mars is a science fiction novel aimed at a younger audience.
+It was written by <entity><mark>Lester del Rey</mark><delete>Andy Weir</delete></entity>
+and published by John C. Winston Co. in 1952, featuring illustrations by Alex Schomburg.
+<contradictory>It ended up having a readership of older boys despite efforts for it to be
+aimed at younger kids .</contradictory>. <invented>The novel inspired the famous Broadway
+musical "Stranded Stars," which won six Tony Awards.</invented> The novel tells a story
+of being stranded on the <entity><mark>Red</mark><delete>Purple</delete></entity> Planet.
+<subjective>I wish the novel had more exciting and thrilling plot twists.</subjective>
+Now detect errors and include edits in the following passage like done in the example above.
+Include error tags <> for ANYTHING YOU CHANGE IN THE ORIGINAL PASSAGE.
+Passage: [PASSAGE_TO_VERIFY]
+Reference: [REFERENCE]
+Edited:
+Table 13: Prompt used for baseline models for detection benchmark.
+We insert the retrieved context at the [Reference:] portion for retrieval-augmented baselines.
+For non-retrieval-augmented baselines, we do not insert [Reference:] component.
+C.3 Details of Human Evaluation on Model Outputs
 We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT.
+Our results are shown in Table  14.
 We anonymize the model results to avoid potential biases during evaluations.
-Annotators count how many of the detection and editing results are correct given extensive web search.
-C.3 Manual Analysis on Generated Data for Fava training
+Annotators count how many of the detection and editing results are correct given extensive web searches.
+Editor avg. $|\mathbf{E}|$ Detect (%) Edit (%)
+Rt+ChatGPT 1.9 23.9 17.2
+Fava (ours) 2.4 55.7 63.5
+Table 14: Human evaluation results. We show the average number of detected errors and the correctness (%) of the fine-grained types and edits.
+C.4 Manual Analysis on Generated Data for Fava training
 We conduct human evaluations on 50 generated data to assess the automatic data creation quality.
@@
 (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed.
-Therefore, we evaluate not only validity but also quality Table  14 shows examples of generated erroneous data from our synthetic data creation pipeline.
+Therefore, we evaluate not only validity but also quality.
+In particular, we evaluate a generated instance from two aspects: (1) validity—whether the model edits violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples.
+We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
+Our analysis revealed that the data generated by our system with iterative insertion obtained an average score of 1.66 for validity assessment and 1.36 for quality assessment.
+Meanwhile, data generated by one-shot prompting scored an average of 1.1 for validity and 0.9 for quality.
+Our human evaluation reveals that our LM-generated edits are indeed of higher quality than the one-shot data creation, and provide realistic factual errors.
+Table  15 shows examples of generated erroneous data from our synthetic data creation pipeline.
 We ask human annotators to score each category for each passage either 0, 1, or 2.
@@
 “Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim.
-<invented>The game won the Shooting Star Award in 2020 for its innovative concept.</invented> Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand.
+¡invented$>$ The game won the Shooting Star Award in 2020 for its innovative concept.$<$/invented$>$ Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand.
 Despite an announced home port for the 3DO Interactive $<$entity$>$ $<$delete$>$ Multiplayer$<$/delete$>$ $<$mark$>$ Singleplayer $<$/mark$>$ $<$/entity$>$ in 1994, the release was ultimately canceled and never reached the market.
@@
 The knapsack problem is a problem in combinatorial optimization where the goal is to determine the optimal selection of items to maximize the total value within a given weight constraint.
-The knapsack problem is a problem in <entity><delete>combinatorial optimization</delete><mark>sequential search</mark></entity> where the goal is to determine the optimal selection of items to maximize the total <entity><delete>value</delete><mark>weight</mark></entity> within a given weight constraint.
+The knapsack problem is a problem in $<$entity$>$ $<$delete$>$ combinatorial optimization$<$/delete$>$ $<$mark$>$ sequential search$<$/mark$>$ $<$/entity$>$ where the goal is to determine the optimal selection of items to maximize the total $<$entity$>$ ¡delete$>$ value¡/delete$>$ ¡mark$>$ weight¡/mark$>$ ¡/entity$>$ within a given weight constraint.
 Table 15: Examples of generated data from automatic data creation.
-C.4 More Ablations
+C.5 More Ablations
 Retrieval ablations.

### v4

@@
 We propose a novel task of automatic fine-grained hallucination detection and construct a new evaluation benchmark, FavaBench, that includes about one thousand fine-grained human judgments on three LM outputs across various domains.
-Our analysis reveals that ChatGPT and Llama2-Chat (70B, 7B) exhibit diverse types of hallucinations in the majority of their outputs in information-seeking scenarios.
-We train Fava, a retrieval-augmented LM by carefully creating synthetic data to detect and correct fine-grained hallucinations.
-On our benchmark, our automatic and human evaluations show that Fava significantly outperforms ChatGPT and GPT-4 on fine-grained hallucination detection, and edits suggested by Fava improve the factuality of LM-generated text.
-[^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/
-Fine-grained Hallucination Detection and Editing for Language Models
-[…]
+Our analysis reveals that ChatGPT and Llama2-Chat (70B, 7B) exhibit diverse types of hallucinations in the majority of their outputs in information-seeking scenarios, highlighting the need to build fine-grained systems.
+To this end, we train Fava, a powerful retrieval-augmented LM by carefully creating synthetic data to detect and correct fine-grained hallucinations.
+Our automatic and human evaluations show that Fava significantly outperforms retrieval-augmented ChatGPT and GPT-4 on fine-grained hallucination detection.
+Furthermore, Fava outperforms widely-used hallucination detection systems on binary detection and shows effectiveness in editing to improve the factuality.
+[^1] Our code, data, and demo are available at https://fine-grained-hallucination.github.io/.
+Figure 1: Overview of our taxonomy, fine-grained hallucination detection task, and Fava.
 1 Introduction
@@
 However, these models are prone to produce factually incorrect or unverifiable statements, often called hallucinations.
-This impedes their deployment in real-world applications for information-seeking contexts Mallen et al.
-(2022).
-Prior work on hallucinations in natural language generation (NLG) often assumes the presence of a specific reference text and focuses on studying faithfulness to the references Ji et al.
-(2023).
-On the contrary, escalating apprehensions have been articulated concerning LM generations that are not grounded in any specific source text, but rather in world knowledge and facts Zhang et al.
-(2023).
-Figure 1: Overview of taxonomy for LM hallucinations and our model, Fava, a fine-grained detection system.
+This impedes their deployment in real-world applications for information-seeking contexts (Mallen et al., 2022; Asai et al., 2024).
+Prior work on hallucinations in natural language generation (NLG) often assumes the presence of a specific reference text and focuses on studying faithfulness to the references (Ji et al., 2023).
+On the contrary, escalating apprehensions have been articulated about LM generations that are not grounded in any specific source text, but rather in world knowledge (Zhang et al., 2023).
 Figure 2: An overview of our fine-grained hallucination taxonomy. We identify 6 fine-grained types representing diverse hallucinations in LM-generated text.
-Several recent work studies automatic hallucination detection Min et al.
-(2023) or editing outputs Gao et al.
-(2022) to address such LM hallucinations.
-These systems typically categorize hallucinations into simplistic binary distinctions like factual or not factual (Figure 1 top).
+Several recent work studies automatic hallucination detection (Min et al., 2023) or editing outputs (Gao et al., 2022) to address such LM hallucinations.
+These systems typically categorize hallucinations into simplistic binary distinctions like factual or not factual (Figure 1 right) or mainly focus on entity-level errors.
 We argue that hallucinations manifest in diverse forms, each requiring varying degrees of careful assessments to verify factuality.
@@
 Conversely, errors involving fabricated entities (e.g., The Messi Diaries in Figure 1) demand thorough verification across multiple sources.
-This underscores the need for a more fine-grained approach to detect hallucination for model development and human verification.
-In this work, we propose automatic fine-grained hallucination detection (Figure 1 bottom), a new task requiring a system to provide precise identification of hallucination sequences, discern different types of hallucinations based on a taxonomy, and suggest refinements.
+This underscores the need for a more fine-grained approach to detect hallucinations.
+In this work, we propose automatic fine-grained hallucination detection (Figure 1 left), a new task requiring a system to provide precise identification of hallucination sequences, discern different types based on a taxonomy, and suggest refinements.
 We focus on hallucinations in information-seeking scenarios, where grounding to world knowledge matters.
-Our taxonomy (Figure 2) hierarchically classifies hallucinations in LM generations into six categories, based on careful pilot studies with NLP experts.
+Our taxonomy (Figure 2) hierarchically classifies hallucinations in LM generations into six categories.
+Our taxonomy is based on iterative pilot studies with NLP experts, to ensure comprehensive coverage of all significant types of hallucinations.
 We construct a new fine-grained hallucination benchmark, FavaBench, by carefully annotating approximately 1,000 responses of three widely used LMs (Llama2-Chat 7B, 70B and ChatGPT [^2] We use gpt-3.5-turbo-0301 throughout this work.) to diverse knowledge-intensive queries.
@@
 Our analysis reveals all models include at least one hallucination in the majority of their responses (e.g., 70.2% in Llama2 7B and 59.8% in ChatGPT).
-In addition to the widely-studied entity-level errors, other error types like unverifiable sentences make up over 60% of LM-generated hallucinations, highlighting the urgent need to study and detect such diverse types.
-We introduce Fava, a new retrieval-augmented LM that can identify and mark hallucinations at the span level using a unified syntax (Figure 1, left bottom).
-Due to the annotation costs, we design an LM-based synthetic data generation process, and train Fava on the 35k resulting instances.
-We compare Fava with state-of-the-art LMs on fine-grained hallucination detection and editing in LM outputs.
+Besides the widely-studied entity-level errors, other types like unverifiable sentences make up over 60% of LM-generated hallucinations, highlighting the urgent need for fine-grained detection.
+Yet, prior methods often only provide statement-level binary hallucination predictions.
+To advance fine-grained hallucination detection, we introduce Fava, a new retrieval-augmented LM that can identify and mark hallucinations at span-level using a unified syntax (Figure 1, left).
+We design an LM-based synthetic data generation and train Fava on the 35k resulting instances.
+We compare Fava with state-of-the-art LMs on fine-grained hallucination detection based on FavaBench.
 Fava significantly outperforms ChatGPT with and without external knowledge by 23.7% on fine-grained hallucination detection.
-Furthermore, Fava effectively corrects hallucinations in diverse LM outputs—Fava editing improves FActScore Min et al.
-(2023) of Alapaca 7, 13B, and ChatGPT by 4.4, 9.3 and 3.3%, respectively.
+On binary hallucination detection, Fava outperforms a widely-used system, FActScore (Min et al., 2023) in addition to other strong baselines leveraging GPT4 or ChatGPT.
+Fava also effectively corrects hallucinations in diverse LM outputs—Fava editing improves a factuality score  (Min et al., 2023) of Alapaca 7, 13B, and ChatGPT by 4.4, 9.3 and 3.3%, respectively.
 2 Related Work
-Type Example ChatGPT Llama2-7b Llama2-70b
+Type Example Chat L-7b L-70b
  Entity Lionel Andrés Messi was born on June 12 24, 1987. 42.7% 48.6% 55.7%
- Relation Lionel Messi acquired was acquired by Paris Saint-Germain. 4.7% 3.4% 2.8%
-  Sentence Messi has yet to be captain for the Argentina football team. 18.9% 15.7% 19.9%
- Invented Messi is known for his famous airplane kick technique. 14.2% 17.4% 9.5%
+ Relation Lionel Messi acquired was acquired by PSG. 4.7% 3.4% 2.8%
+  Sentence Messi was never captain for Argentina football team. 18.9% 15.7% 19.9%
+ Invented Messi is known for his famous airplane kick. 14.2% 17.4% 9.5%
  Subjective Lionel Messi is the best soccer player in the world. 9.9% 8.6% 5.8%
- Unverifiable In his free time, Messi enjoys singing songs for his family. 9.6% 6.3% 6.3%
-Table 1: Distribution of different errors across ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs and examples of hallucinations for each type from our human-annotated data
+ Unverifiable When free, Messi enjoys singing songs for his family. 9.6% 6.3% 6.3%
+Table 1: Distribution of different errors across ChatGPT (Chat), Llama2-Chat-7B (L-7B) and Llama2-Chat-70B (L-70B) outputs and examples of hallucinations for each type.
 Hallucinations in NLG.
-Prior taxonomies proposed for summarization Pagnoni et al.
-(2021) or text simplifications Devaraj et al.
-(2022) often assume the existence of specific source text and focus on faithfulness to the source Ji et al.
-(2023).
+Prior taxonomies proposed for summarization (Pagnoni et al., 2021), text simplifications (Devaraj et al., 2022) or knowledge-grounded dialogue (Dziri et al., 2022) often assume the existence of specific source text and focus on faithfulness to the source (Ji et al., 2023).
 As this is qualitatively different from LMs’ factual hallucinations grounded in world knowledge, we extend prior work and build a new taxonomy for hallucinations in LM outputs.
 Detecting and editing hallucinations in LMs.
-Several recent or concurrent studies propose methods to predict factuality by giving binary labels of a statement being factual or not Manakul et al.
-(2023); Min et al.
-(2023), or focusing on entity-level hallucination editing Gao et al.
-(2022); Chen et al.
-(2023).
+Several recent or concurrent studies propose methods to predict factuality by giving binary labels of a statement being factual or not (Manakul et al., 2023; Min et al., 2023), focusing on entity-level hallucination editing (Gao et al., 2022; Chen et al., 2023), or concentrating purely on detection (Li et al., 2024).
 Recent surveys have attempted to categorize hallucinations in LM generations, often stretching the definition of hallucinations to broad error types in LM-generated text, which cannot be traced to specific spans in text (Huang et al., 2023; Zhang et al., 2023; Rawte et al., 2023).
 We propose a fine-grained taxonomy for hallucinations in long-form text generation for information-seeking context to improve factual consistency.
-While prior work often develops binary factuality verification systems on top of proprietary LMs Chern et al.
-(2023), we train a fine-grained detection system using our carefully designed data generation pipeline.
+While prior work often develops binary factuality verification systems on top of proprietary LMs (Chern et al., 2023), we train a fine-grained detection system using our carefully designed data generation pipeline.
 Fact verification for human-written claims.
 While research on hallucination detection focuses on identifying errors in model-generated text, a related area of research focuses on identifying factual inaccuracies in human written claims (Bekoulis et al., 2021).
-Many large-scale datasets have been proposed for fact verification given Wikipedia Thorne et al.
-(2018), scientific articles (Wadden et al., 2020) or news documents (Wang, 2017).
-Several systems have been developed for fact-checking in these settings (Thorne and Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text.
+Many large-scale datasets have been proposed for fact verification given Wikipedia (Thorne et al., 2018), scientific articles (Wadden et al., 2020) or news documents (Wang, 2017).
+Several systems have been developed for fact-checking in these settings (Thorne & Vlachos, 2018; Schuster et al., 2021; Nakov et al., 2021), but have not been tested for LM-generated text.
 While our work focuses on hallucination detection in LM generations, Fava can be adapted to fact-check human written claims.
@@
 We conducted a pilot annotation with nine NLP experts to ensure good coverage across diverse error types.
-Figure 2 shows our taxonomies to classify LM hallucinations and Table 1 shows examples of each type of hallucination.
+Figure 2 shows our taxonomies to classify LM hallucinations.
+Table 1 shows examples of each type of hallucination.
 We group hallucinations into two major categories: statements that contradict world knowledge (Type 1) and unverifiable statements (Types 2, 3, and 4).
-(1a)  Entity : Contradictory entity errors are a sub-category within contradictory statement errors where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
+(1a)  Entity : Contradictory entity errors are a sub-category within Type 1, where an entity in a statement is incorrect and changing that single entity can make the entire sentence factually correct.
 (1b)  Relation : Contradictory relation errors are another sub-category within contradictory statements where a semantic relationship (e.g., verbs, prepositions, or adjectives) in a statement is incorrect.
-(1c)   Sentence : Contradictory sentence errors refer to cases where a full statement entirely contradicts relevant evidence from the web.
+(1c)   Sentence : Contradictory sentence errors refer to cases where a full statement entirely contradicts relevant evidence from the web, and cannot be solved via phrase-level edits.
 (2)  Invented : Invented errors refer to statements where the LM generates an entirely fabricated entity that doesn’t exist based on world knowledge.
 Fictional entities in creative work aren’t included.
-(3)  Subjective : Subjective errors refer to an expression that lacks universal validity and is often influenced by personal beliefs, opinions, or biases.
-(4)  Unverifiable : This refers to statements where the LM output contains facts, but none of the retrieved evidence from the web can directly support or contradict the fact (e.g., personal matters).
+(3)  Subjective : Subjective errors refer to expressions about existing entities that lack universal validity.
+These statements often do not contain facts and are influenced by personal beliefs or opinions.
+(4)  Unverifiable : These are statements where the LM output contains facts, but no retrieved evidence from the web can directly support or contradict the fact (e.g., private details).
+[^4] Differing from  Subjective errors which focus on opinionated statements lacking facts,  Unverifiable specifically focuses on fact-based statements that just lack evidence.
 [^5] While  Invented and  Unverifiable are highly related,  Invented are the hallucinations where we can verify that some core entities or subjects of the sentences don’t exist.
@@
 We evaluate systems’ abilities concerning identifying fine-grained error types in the model-generated text $e^{type}$ (Task 1; Figure 1 < Entity >) and editing factual errors $e^{text}$ accordingly (Task 2; Figure 1 Barcelona $\rightarrow$ Miami).
-Figure 3: Distribution of hallucination types in ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs across four datasets of diverse information-seeking queries based on our human annotated benchmark.
+Figure 3: Distribution of hallucination types in ChatGPT, Llama2-Chat-7B and Llama2-Chat-70B outputs across four datasets of diverse information-seeking queries.
 Task 1: Fine-grained hallucination detection.
@@
 Given an output $y$ consisting of $L$ sentences, we assume the availability of ground-truth error type annotations $e^{*t}_{i}\in\{\texttt{TRUE},\texttt{FALSE}\}$, which is a binary label of an error type $t$ existing in the $i$th sentence (TRUE) or not (FALSE).
-Following the fact verification literature Thorne et al.
-(2018); Schuster et al.
-(2021); Feng et al.
-(2023) of computing precision and recall, we measure those metrics while we extend them for each error type.
-For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as follows:
-$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=\texttt{TRUE}]}$ (1)
-$\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}=\texttt{TRUE}]}$ (2)
+Following the fact verification literature (Thorne et al., 2018; Schuster et al., 2021; Feng et al., 2023) of computing precision and recall, we measure those metrics while we extend them for each error type.
+For each type, a system predicts $e^{t}_{i}$ and we evaluate precision and recall as:
+$\text{Prec}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=\texttt{TRUE}]}{\rm,~{}~{}~{}}\text{Recall}^{t}=\frac{\sum_{i\in L}\mathbbm{1}[e^{t}_{i}=e^{*t}_{i}]}{\sum_{i\in L}\mathbbm{1}[e^{*t}_{i}=\texttt{TRUE}]}$ (1)
 Precision indicates the proportion of how many of the model’s predictions of an error type $t$ existing in the $i$th sentence is correct, while recall indicates how many of the error sentence TRUE is identified by the model.
@@
 Fine-grained detection can be simplified into a binary classification task, in which a system predicts if a sentence $s_{i}$ includes any factual errors or not.
-[^5] This is similar to FActScore Min et al.
-(2023), while we predict and evaluate factuality at the sentence level without extracting atomic facts.
+[^6] This is similar to FActScore (Min et al., 2023), while we predict and evaluate factuality at the sentence level without extracting atomic facts.
+We compare Fava against FActScore on this task.
 Task 2: Hallucination editing.
@@
 In the editing task, the system is expected to suggest fine-grained edits to improve the factuality of the LM output.
-We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore Min et al.
-(2023), as: $f(\hat{y})-f(y)$, where $f$ indicates an estimated factuality scores and $\hat{y}$ indicates edited output.
+We evaluate a system’s ability to improve the factuality of given output $y$, by comparing scores estimated by off-the-shelf systems, such as FActScore (Min et al., 2023), as: $f(\hat{y})-f(y)$, where $f$ indicates an estimated factuality scores and $\hat{y}$ indicates edited output.
 4 Benchmark: FavaBench
-We create FavaBench to facilitate model development and evaluation in fine-grained hallucination detection and understand how prevalent those different types of hallucinations are.
-FavaBench consists of multi-way fine-grained annotations on three LM responses to queries in multiple domains.
-Source prompts.
-The source prompts include a collection of 200 information-seeking queries, spanning four different data sources. See examples in Appendix Table 4.
-  • Knowledge-intensive queries sampled from the Open Assistant dataset Köpf et al. (2023), by prompting GPT-4 to judge whether each query from the dataset requires world knowledge.
-  • A sample of Open QA prompts from the No Robots dataset Rajani et al. (2023).
-  • Instruction-following prompts requiring more reasoning and knowledge curated by the authors.
-  • Synthetically-created prompts 50 prompts that require more fine-grained knowledge, by converting data-to-text WebNLG dataset Gardent et al. (2017) using a template.
+FavaBench consists of around 1k fine-grained annotations on three LM responses to queries in multiple domains.
+Source prompts. The source prompts include a collection of 200 information-seeking queries, spanning four different data sources. See examples in Appendix Table 6.
+  • 50 Knowledge-intensive queries from Open Assistant (Köpf et al., 2023), by prompting GPT-4 to judge whether each query from the dataset requires world knowledge.
+  • 50 Open QA prompts from the No Robots dataset (Rajani et al., 2023).
+  • 50 instructions requiring more reasoning and knowledge by the authors.
+  • 50 synthetically-created prompts that require more fine-grained knowledge, by converting data-to-text WebNLG dataset (Gardent et al., 2017) using a template.
 Annotation details and quality.
 We obtain responses to the collected source prompts using ChatGPT (gpt-3.5-turbo-0301;  Ouyang et al.
-2022) Llama2-Chat 7B and Llama2-Chat 70B Touvron et al.
-(2023) in a zero-shot manner and collect 600 responses to our diverse information-seeking prompts.
+2022) Llama2-Chat 7B and Llama2-Chat 70B (Touvron et al., 2023) in a zero-shot manner and collect 600 responses to our diverse information-seeking prompts.
 We recruited 20 students (ten undergraduate and ten NLP graduate students) to annotate the factual accuracy of the responses based on our proposed taxonomy.
 Each instance is annotated by two annotators who completed 45-minute in-person and virtual training sessions.
-[^6] Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for each annotation.
-Our annotation interface and details are in Appendix A.
-To validate our annotation quality, we calculated inter-annotator agreement using Cohen kappa scores and found that our annotators had high agreement in detection across passages, with 75.1% agreement in detection at the sentence level and 60.3% agreement in exact error type detection at the sentence level across passages.
-Appendix Table 6 shows examples of our annotation results.
+[^7] Annotating each response takes approximately 10 minutes, and we pay USD 3.5 for annotation.
+We provide annotators with detailed instructions, examples, and training to ensure high-quality annotations.
+Our annotation interface and details are in Appendix A and examples in Appendix Table 8.
+To validate our annotation quality, we calculated inter-annotator agreement using Cohen kappa scores.
+Our annotators have high agreement in detection across passages, with 75.1% agreement in detection at the sentence level and 60.3% agreement in exact error type detection.
 Analysis on annotated data.
@@
 Invented  are more common in Llama2-Chat models.
-Fewer errors in No Robots subsets than other subsets may be because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How long was the Revolutionary War?), which is often memorized by LMs Mallen et al.
-(2022).
+Fewer errors in No Robots subsets than other subsets may be because their seed prompts are less knowledge-intensive, or ask about popular factual knowledge (e.g., How long was the Revolutionary War?), which is often memorized by LMs (Mallen et al., 2022).
 Figure 4: Overview of high-quality synthetic data generation process in Fava.
@@
 Fava is trained on high-quality synthetic training data to identify hallucinations, incorporating retrieved knowledge.
-Overview.
 Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$.
@@
 $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1.
-While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2).
-Reliance on such API models also hurt reproducibility.
-Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.1) and fine-tune a smaller yet powerful expert LM on the generated data (Section 5.2).
+While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (Table 2).
+Reliance on such models also hurts reproducibility.
+Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.1) and fine-tune a smaller yet powerful LM on the generated data (Section 5.2).
 5.1 Synthetic Training Data Curation
-To train our $\mathcal{M}_{edit}$ model, we require a large number of training instances, each of which includes $(c,y,y^{*})$, where $c$ is the gold context (original Wikipedia paragraph), $y$ is an erroneous LM output, and $y^{*}$ is an output with error tags and correct editing based on $c$.
-Fava leans to take $(c,y)$ and generate $y^{*}$.
-Yet, such data is not publicly available, and our annotation is expensive to scale up.
-Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al.
-(2022); Wang et al.
-(2023b); Asai et al.
-(2023), we introduce a new data creation method tailored to our fine-grained hallucination taxonomies.
-Specifically, our data creation pipeline consists of three steps: Step 1: seed passage generations that generates a large, diverse collection of seed passages $c$, Step 2: error insertions which inserts factually incorrect statements informed by our our hallucination taxonomy to generate erronous passages $y$, and Step 3: post-processing which prunes and curate the final training examples.
-Figure 4 shows the overview of the pipeline.
+To train our $\mathcal{M}_{edit}$ model, we require a large number of training instances, each of which includes $(c,y,y^{*})$, where $c$ is the gold context, $y$ is an erroneous LM output, and $y^{*}$ is an output with error tags and correct editing based on $c$.
+Fava learns to take $(c,y)$ and generate $y^{*}$.
+Inspired by prior work that leverages LMs to generate synthetic training data (Balachandran et al., 2022; Wang et al., 2023b; Asai et al., 2023), we introduce a new data creation method tailored to our fine-grained hallucination taxonomies.
+Specifically, our data creation pipeline (refer Figure 4) consists of three steps: Step 1: seed passage generations that generates a large, diverse collection of seed passages $c$, Step 2: error insertions which inserts factually incorrect statements to generate erronous passages $y$, and Step 3: post-processing which curates the final training examples.
 Seed passage generation.
@@
 Later on, errors are inserted into these seed passages to generate erroneous text, $y$.
-We base our seed passage collection on Wikipedia articles and available Question Answering (QA) datasets Kwiatkowski et al.
-(2019) since they are both good sources to evaluate factuality.
-However, the text genre is quite limited.
-To diversify the passage genres, we paraphrase Wikipedia passages into different genres to enable Fava to be adaptable and effective across various textual formats.
-More specifically, we randomly sample 35,074 Wikipedia articles $c$ and use them as our gold reference passages.
+We base our seed passage collection on Wikipedia articles and sampled 5k QA pairs from Natural Questions (Kwiatkowski et al., 2019).
+We paraphrase Wikipedia passages into different genres to enable Fava to be adaptable and effective across various textual formats.
+Specifically, we randomly sample 35,074 Wikipedia articles $c$ and use them as our gold reference passages.
 We then randomly sample one of the pre-specified genre types (e.g.
 blog article, tweet, etc.) for each passage and use ChatGPT to paraphrase the original passage into the given genre.
-In addition, we use 5,000 examples from QA datasets to further diversify the passage collections.
-In particular, we prompt ChatGPT with two QA demonstrations to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles.
-See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 7.
+[^8] See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 9.
 Error insertion.
@@
 To control error distributions in the training data and incorporate our hallucination taxonomy, we prompt LMs to insert errors from our taxonomy one by one.
-In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type).
-We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types.
-Therefore, we use ChatGPT and GPT-4 interchangeably for six different types.
+We use ChatGPT and GPT-4 interchangeably for six different types.
 [^9] We use ChatGPT (gpt-3.5-turbo-0301) for the Type 1 and Type 3 while using GPT-4 for other types.
-Given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retaining previously inserted errors and original text.
-The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution.
-After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
+See detailed discussions in Appendix B.
+Given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retaining previously inserted errors.
+Models mark phrases or sentences for deletion along with their error type and insert phrases and sentences with insertion tags.
+Now we have an erroneous text $y$, which inserts multiple errors into the original text $t$.
 Post-processing.
 We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$.
-We filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types).
-We also retrieve additional four relevant Wikipedia paragraphs using Contriever-MSMARCO Izacard et al.
-(2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$.
-Statistics of data.
+We filter out the examples violating our editing rules (e.g., inserting errors without marking them up with tags and error types).
+We also retrieve additional four relevant Wikipedia paragraphs using Contriever-MSMARCO (Izacard et al., 2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$.
+Statistics of data and human evaluation.
 After analyzing ablations with varying training instance sizes (Section 7.2), we generated 35,074 training instances,  [^10] While we see performance improvements up to 30k, the API costs for training data generation were already \$3k, reaching our initial budget.
@@
 All of the error types are almost equally distributed ($\sim$15 % each) with 3.1 errors inserted for Wikipedia subsets, and 1.4 errors inserted for the QA pairs on average.
-We present the error distributions in Appendix Table 8 and a detailed human evaluation in Appendix C.4.
+We present the error distributions in Appendix Table 10.
+We conduct human evaluations on randomly sampled 50 examples and find that the generated data meets the expected criterion in most cases.
+See details in Appendix C.4.
 5.2 Training and Inference
@@
 We empirically found initializing Fava with this model gives slightly better performance than the pre-trained Llama2.
-At inference time, we retrieve the top five documents from Wikipedia, [^9] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder.
+At inference time, we retrieve the top five documents from Wikipedia, [^11] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever.
 using Contriever-MSMARCO, and input them together with an LM output that may include hallucinations.
@@
 6 Experiments
-Generator: ChatGPT Generator: Llama2-Chat 70B
-Editor  ent  rel  con  inv  subj  unv OA Bi  ent  rel  con  inv  subj  unv OA Bi
-ChatGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
-Rt+ChatGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
-GPT4 38.6 16.6 17.9 22.2 50.0 17.2 34.2 60.8 55.5 60.0 21.2 15.4 2.0 25.0 42.5 74.2
-Fava (ours) 54.5 25.0 66.7 16.7 70.5 35.3 48.1 79.6 57.3 34.5 27.7 52.2 31.25 43.4 47.2 80.3
-Table 2: Fine-grained detection F1. OA and Bi indicates overall and binary predictions.
 6.1 Experiments for Hallucination Detection
@@
 We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.2.
+We evaluate both fine-grained detection and binary detection settings.
+Note that fine-grained hallucination detection is our new proposed task, as prior work mostly focuses on the binary detection setting.
 Baselines.
-For fine-grained hallucination detection tasks, we test three baselines.
-ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations.
-GPT4 (gpt-4) does the exact same but prompts GPT4 instead.We were unable to include GPT4 with retrieval in our baseline due to high costs.
+Fine-grained hallucination is a new task and we test three baselines that use state-of-the-art proprietary LMs: ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations.
+GPT4 (gpt-4) does the exact same but prompts GPT4 instead.
+We were unable to include GPT4 with retrieval in our baseline due to high costs.
 Since Fava uses retrieval augmentation, for comparability we include Rt+ChatGPT.
 This baseline prompts ChatGPT using the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt.
-[^11] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
+[^13] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or violate formats.
+We additionally evaluate FActScore (Min et al., 2023) for the binary detection task only.
+This model-based metric prompts ChatGPT and GPT 3 (davinci-003) to decompose a response into a set of atomic facts and verify factuality for each using passages from a designated Wikipedia article.
 6.2 Experiments for Hallucination Editing
 Evaluation data.
-For editing evaluations, we use open-ended text generation data and evaluate models’ factuality based on a metric designed to measure the factuality of text, FActScore Min et al.
-(2023).
-This model-based metric prompts ChatGPT and GPT 3 (davinci-003) to decompose a response into a set of atomic facts and verify factuality for each using passages from a designated Wikipedia article.
-We use the original biography generation task proposed with FActScore.
+For editing evaluations, we use open-ended text-generation data and evaluate models’ factuality based on a metric designed to measure the factuality for the target task.
+In particular, we use the biography generation task proposed by FActScore (Min et al., 2023) to evaluate the effectiveness of editing to reduce hallucinations based on edited outputs’ FActScore results.
 Baselines.
@@
 We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents.
-This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at the individual detection level.
-Due to the cost of evaluating the factuality of all verification-worthy statements Min et al.
-(2023), we do not evaluate recall.
+This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at each detection level.
 7 Results and Analysis
+Generator: ChatGPT Generator: Llama2-Chat 70B
+Editor  ent  rel  con  inv  subj  unv Avg.  ent  rel  con  inv  subj  unv Avg.
+ChatGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 24.7 15.6 26.7 11.0 17.6 12.8 24.1
+Rt+ChatGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 33.7 24.2 24.0 22.2 17.8 4.7 27.8
+GPT4 38.6 16.6 17.9 22.2 50.0 17.2 34.2 55.5 60.0 21.2 15.4 2.0 25.0 42.5
+Fava (ours) 54.5 25.0 66.7 16.7 70.5 35.3 48.1 57.3 34.5 27.7 52.2 31.3 43.4 47.2
+Table 2: Results on fine-grained detection (metric: F1). Full results are in Appendix.
 7.1 Results
-Detection results.
+Generator
+Model Chat Llama
+Chat 50.1 68.4
+R+Chat 64.8 72.8
+GPT4 60.8 74.2
+FAct 67.7 71.2
+Fava 79.6 80.3
+Table 3: Results of binary detection.
+Generator
+Editor CGPT Al-13B Al-7B
+No Edit 66.7 42.5 38.8
+[…]
+Rt-CGPT 62.7 -4.0 43.9 +1.4 39.2 +0.4
+Llama 52.6 -14.1 22.7 -19.8 18.6 -20.2
+Rt-Llama 58.7 -8.0 48.6 +6.1 32.2 -6.6
+[…]
+Table 4: Results of editing.
+Settings Score
+No Edit 42.5
+(Top 1) 44.2 +1.7
+(Top 5) 47.0 +4.5
+Reranked 47.7 +5.2
+Entity 50.1 +7.6
+Table 5: Ablations of
+retrieval for editing.
+Fing-grained detection results.
 Table 2 shows the fine-grained detection accuracy, namely F1 scores, and binary prediction F1 of Fava and baselines.
-We provide the full precision and recall scores in Appendix Tables 11 and 12, respectively.
+We provide the full precision and recall scores in Appendix Tables 13 and 14, respectively.
 Fava significantly outperforms ChatGPT, Ret-ChatGPT, or GPT-4 on both fine-grained error detection and binary error detection.
@@
 Those two error types often require intensive search over many documents beyond the top few ones, while Fava by default only considers the top five documents.
-Generator
-Editor ChatGPT Al-13B Al-7B
-No Edit 66.7 42.5 38.8
-[…]
-Rt-ChatGPT 62.7 -4.0 43.9 +1.4 39.2 +0.4
-Llama 52.6 -14.1 22.7 -19.8 18.6 -20.2
-Rt-Llama 58.7 -8.0 48.6 +6.1 32.2 -6.6
-[…]
-Table 3: Overall Editing results on FActScore Min et al. (2023). Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
+Table 7.1 shows the binary hallucination detection performance of Fava, baselines models, and FActScore (FAct).
+Fava outperforms all other models in detecting hallucinations by a large margin.
 Editing results.
-Table 3 shows the results of the editing task on FActScore.
+Table 7.1 shows the results of the editing task on the biography generation task.
 Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model in size.
-We found that ChatGPT is often confused with different errors and has a tendency to over-predict contradictory or subjective statements.
-Fava shows the largest gains in FActScore, showing the effectiveness of editing, in addition to error type detection.
+We found that ChatGPT is often confused with different errors and tends to over-predict contradictory or subjective statements.
+Fava shows the largest gains in FActScore, showing the effectiveness of editing and error type detection.
 Human evaluation results.
@@
 We assess our model on varying sizes of synthetic training data (10k, 20k, and 30k instances) and examine their performance on fine-grained hallucination detection.
-Figure 4(b) demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
+Appendix Figure 7 demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
 Effect of retrieval.
 We assess Fava’s editing performance on Alpaca 13B outputs with varying retrieved evidence: top 1 document; top 5 documents; reordering the top 5 documents; top 4 documents and mix them with an introductory paragraph of the target entity (entity matching).
-Table 4(a) shows the experimental results, and more detailed results are available in Appendix Table  16.
+Table 5 shows the experimental results.
+More detailed results are in Appendix Table  18.
 We found that retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%.
-Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effect of context ordering, as reported by prior work Liu et al.
-(2023).
+Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effect of context order, as reported by prior work (Liu et al., 2023).
 Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations.
 These results indicate although Fava demonstrates strong capabilities, there’s room for improvement in retrieving and incorporating many references.
-Settings Score
-No Edit 42.5
-(Top 1) 44.2 +1.7
-(Top 5) 47.0 +4.5
-Reranked 47.7 +5.2
-Entity 50.1 +7.6
-(a) Retrieval
-(b) Training Size
-Figure 5: (a): FActScore results with different retrieval settings. (b): Detection scores across different training instance values.
 8 Conclusions
@@
 Empirical results show that Fava significantly outperforms strong baselines by a large margin on both editing and fine-grained detection tasks, while still large room for improvements for automated fine-grained error detection and editing.
-Limitations
-This work focuses on hallucinations in LM generations in an information-seeking context, where factual correct matters, and we build a new taxonomy, benchmark, and model for this.
-Errors in other contexts (e.g., commonsense reasoning, numerical reasoning) are beyond our focus.
-Our training data uses Wikipedia passages as gold context, and Fava uses Wikipedia passages as references at inference time.
-Yet, Wikipedia may not cover all of the necessary information to verify the factuality of LM-generated statements Asai and Choi (2021), and it is an open question how well Fava performs when it uses other types of documents as reference documents.
-Our analysis and findings have been based only on the English language and our scope is limited to information-seeking scenarios.
-Our model, Fava, and benchmark, FavaBench, were also trained on English text data.
-To understand fine-grained hallucinations in diverse languages or scenarios, more analysis and experiments need to be done.
-Our annotations were collected with undergraduate and graduate computer science students who attended 45-minute training sessions to guarantee quality.
-Furthermore, annotations were reviewed for quality by us authors and annotator agreement was calculated to ensure understanding within annotators.
-Yet, given the complexity of our annotation tasks, there may be subjectivity or errors in annotations.
-We will release our data publicly to enable further investigations.
-It is difficult to guarantee that our proposed hallucination taxonomy is all-encompassing.
-To address this concern, we asked annotators to include hallucination types not found in our taxonomy during the annotation process.
-We received no additions to our taxonomy during this, indicating our taxonomy is thorough.
-We evaluate Fava to determine feasibility and quality. Although this doesn’t assure quality outputs from Fava consistently, the positive results indicate quality within our defined scope and setup.
-Ethics Statement and Broader Impacts
-LM-generated text is prone to hallucinations, which can result in severe real-world risks such as spreading misinformation.
-In this work, we attempt to understand such prevalent hallucinations by providing a new taxonomy, a human-annotated benchmark, and a new expert LM trained for fine-grained hallucinations.
-We hope our work will inspire future work and encourage more work focusing on fine-grained hallucination detections.
-Yet, our resulting models are not perfect and may overlook hallucinations in LM outputs and we acknowledge such limitations in our paper.
-For evaluating the detection capabilities of our model, we use our annotation data.
-Our annotation collection process did not involve gathering private information or data from our annotators.
-Additionally, of the four datasets we used for annotation prompts, three are publicly available datasets containing no personal information.
-The remaining prompt dataset was curated by us authors and does not include personal information.
-Our annotators were all paid for their time in completing annotations at the rate of \$20 per hour for 10 hours of work across a week.
-They all disclosed the payment amount and estimated time before beginning annotations.
-None of their data was collected while constructing the annotation dataset.
-For evaluating the editing capabilities of our model, we use FActScore Min et al.
-(2023) which has publicly available code and data and is based on Wikipedia.
 Acknowledgement
-We thank Sewon Min for fruitful discussions in the early stages.
-Akari Asai is supported by the IBM Fellowship.
+We thank Sewon Min for fruitful discussions in the early stages and UW NLP members and annonymous reviewers for their insightful feedback on the draft.
 We thank Stability AI for providing computing to train and evaluate the LMs in this work, and Microsoft Accelerate Foundation Models Research Program for the access to Azure and OpenAI APIs.
@@
 This work was funded in part by the DARPA MCS program through NIWC Pacific (N66001-19-2-4031), NSF IIS-2044660, and gifts from AI2.
+We gratefully acknowledge support from NSF CAREER Grant No.
+IIS2142739, NSF Grants No.
+IIS2125201, IIS2203097, and gift funding from Google, MSR, and OpenAI.
 Appendix
@@
 Table 7: Hallucination percentages in our human-annotated data.
-Dataset
-Annotated Passage Example
-WebNLG
-A Long Long Way is a novel written by Sebastian Barry.
+Dataset Annotated Passage Example
+WebNLG A Long Long Way is a novel written by Sebastian Barry.
 It was first published in 2005 by $<$entity$>$Faber and Faber viking press$<$/entity$>$ , a $<$entity$>$ British American$<$ /entity$>$ publishing company.
@@
 The novel was critically acclaimed and received several literary awards, $<$contradictory$>$ including the 2005 Dublin Literary Award $<$/contradictory$>$ and was shortlisted for the Man Booker Prize in the same year.
-OpenAssistant
-$<$subjective$>$ The most luxurious airline flying within the continental United States is generally regarded as JetBlue Airways$<$/subjective$>$ , particularly for its Mint business class service.
+OpenAssistant $<$subjective$>$ The most luxurious airline flying within the continental United States is generally regarded as JetBlue Airways$<$/subjective$>$ , particularly for its Mint business class service.
 Mint offers a private suite experience, featuring fully lie-flat seats, ample legroom and storage, premium food and beverage options, access to onboard WiFi and live TV streaming, and turn-down service complete with duvets and pillows.
-Instruction Following
-Some of Afghanistan’s secondary cricket home grounds include:
+Instruction Following Some of Afghanistan’s secondary cricket home grounds include:
 1.$<$contradictory$>$Kandahar Cricket Stadium, Kandahar$<$/contradictory$>$
@@
 Appendix B More Details of Synthetic Data Creation and Training
+Quality of error insertions.
+In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type).
+We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types.
 Text diversification results.
@@
 Seed Passage Diversification Examples
-Genre Topic
-Output
-Article Community Reinvestment Act
-The Community Reinvestment Act (CRA) is a federal law that encourages commercial banks and savings associations to help meet the needs of borrowers in low- and moderate-income neighborhoods.
+Genre Topic Output
+Article Community Reinvestment Act The Community Reinvestment Act (CRA) is a federal law that encourages commercial banks and savings associations to help meet the needs of borrowers in low- and moderate-income neighborhoods.
 The Act was passed in 1977 by Congress to reduce discriminatory credit practices against low-income neighborhoods.
-Essay National Dodgeball League
-The National Dodgeball League is the only professional dodgeball league in the US, and was founded in 2004.
+Essay National Dodgeball League The National Dodgeball League is the only professional dodgeball league in the US, and was founded in 2004.
 It is headed by Commissioner Edward Prentiss out of Hopkins, MN and is composed of 24 professional teams that are divided into the National and American Dodgeball Conferences.
 In the past, the tournament has featured teams from Canada, Denmark, UK, Japan, New Zealand and Australia.
-Blog Post Red Channels
-In the beginning of the 1950s, the ”Counterattack” journal published a document, ”Red Channels,” which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association.
+Blog Post Red Channels In the beginning of the 1950s, the ”Counterattack” journal published a document, ”Red Channels,” which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association.
 Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history.
 ”Red Channels” effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
-Social Media Post Sandman
-Check out The Sandman - a semi-retired American professional wrestler!
+Social Media Post Sandman Check out The Sandman - a semi-retired American professional wrestler!
 He is best known for his career with Extreme Championship Wrestling (ECW), where he held the ECW World Heavyweight Championship a record five times.
@@
 #Wrestling #ECW #WCW #TNA #WWE
-News Report I Don’t Mind (James Brown song)
-James Brown’s ”I Don’t Mind” has made its way up the R&B ”Billboard” charts, hitting number four.
+News Report I Don’t Mind (James Brown song) James Brown’s ”I Don’t Mind” has made its way up the R&B ”Billboard” charts, hitting number four.
 The track was released in 1961 and also reached number 47 on the Pop ”Billboard” charts.
 Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
-Summary Robert Lui
-Robert Lui was born on February 23, 1990, and is a professional rugby league footballer from Australia.
+Summary Robert Lui Robert Lui was born on February 23, 1990, and is a professional rugby league footballer from Australia.
 He plays either as a halfback or five-eighth for the Townsville Blackhawks in the Queensland Cup.
-Speech George Sperling
-Dear fellow Americans, I am here to propose a solution to improve American Sign Language communication.
+Speech George Sperling Dear fellow Americans, I am here to propose a solution to improve American Sign Language communication.
 George Sperling suggests that with a sevenfold reduction in the bandwidth for video transmission, we can achieve this.
@@
 Let us not forget our roots and make a change for the better.
-Presentation Intro Blissful Ignorance Effect
-Have you ever wondered why sometimes people who know less about a product seem to enjoy it more than those who have researched it thoroughly?
+Presentation Intro Blissful Ignorance Effect Have you ever wondered why sometimes people who know less about a product seem to enjoy it more than those who have researched it thoroughly?
 This phenomenon is called the Blissful Ignorance Effect and it’s a fascinating topic in consumer behavior studies.
 Our presentation today will explore this effect and why it happens.
-Brochure Nidulariaceae Fungi
-Explore the fascinating world of Nidulariaceae fungi!
+Brochure Nidulariaceae Fungi Explore the fascinating world of Nidulariaceae fungi!
 This family, found in most ecological regions, includes five different genera, each with its own unique characteristics.
 With their tiny egg-filled structures, these fungi are a wonder to behold!
-Text Message Ashes and Diamonds
-Just found out about this book called Ashes and Diamonds by Jerzy Andrzejewski. It’s set during the last few days of WWII. The main character, Maciek, has to kill a Communist soldier. Sounds intense!
+Text Message Ashes and Diamonds Just found out about this book called Ashes and Diamonds by Jerzy Andrzejewski.
+It’s set during the last few days of WWII.
+The main character, Maciek, has to kill a Communist soldier.
+Sounds intense!
 Table 9: Text diversification prompts.
@@
 Our base model is Llama2-Chat 7b trained using 4xA40 GPUs.
-Our training code is based off Open-Instruct Wang et al.
-(2023a) [^13] https://github.com/allenai/open-instruct/blob/main/scripts/finetune_with_accelerate.sh.
+Our training code is based off Open-Instruct (Wang et al., 2023a) [^15] https://github.com/allenai/open-instruct/blob/main/scripts/finetune_with_accelerate.sh.
 Table 11 shows the training hyperparameters.
@@
 Llama2-Chat 7B F1 results
-Table  10 reports the detection results for the Llama2-Chat 7B generations on our curated datasets.
+Table  12 reports the detection results for the Llama2-Chat 7B generations on our curated datasets.
 Generator: Llama2-Chat 7B
@@
 Overall precision and recall.
-Tables 11 and 12 report the precision and recall on our curated datasets.
+Tables 13 and 14 report the precision and recall on our curated datasets.
 ChatGPT Generations Llama2-Chat 70B Generations
@@
 C.2 Benchmark Prompts
-Figure C.2 shows the prompt used for baseline models when evaluating on the detection benchmark.
+Table 15 shows the prompt used for baseline models when evaluating on the detection benchmark.
 Given a passage with factual errors, identify any <entity>, <relation>,
@@
 We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT.
-Our results are shown in Table  14.
+Our results are shown in Table  16.
 We anonymize the model results to avoid potential biases during evaluations.
@@
 In particular, we evaluate a generated instance from two aspects: (1) validity—whether the model edits violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples.
-We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
+We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better).
 Our analysis revealed that the data generated by our system with iterative insertion obtained an average score of 1.66 for validity assessment and 1.36 for quality assessment.
@@
 Our human evaluation reveals that our LM-generated edits are indeed of higher quality than the one-shot data creation, and provide realistic factual errors.
-Table  15 shows examples of generated erroneous data from our synthetic data creation pipeline.
+Table  17 shows examples of generated erroneous data from our synthetic data creation pipeline.
 We ask human annotators to score each category for each passage either 0, 1, or 2.
@@
 As a comparison, we also generate training data with one-shot generation.
-Original Passage
-Automatic Data Generations for Training
+Original Passage Automatic Data Generations for Training
 “Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim.
@@
 The knapsack problem is a problem in combinatorial optimization where the goal is to determine the optimal selection of items to maximize the total value within a given weight constraint.
-The knapsack problem is a problem in $<$entity$>$ $<$delete$>$ combinatorial optimization$<$/delete$>$ $<$mark$>$ sequential search$<$/mark$>$ $<$/entity$>$ where the goal is to determine the optimal selection of items to maximize the total $<$entity$>$ ¡delete$>$ value¡/delete$>$ ¡mark$>$ weight¡/mark$>$ ¡/entity$>$ within a given weight constraint.
+The knapsack problem is a problem in $<$entity$>$ $<$delete$>$ combinatorial optimization$<$/delete$>$ $<$mark$>$ sequential search$<$/mark$>$ $<$/entity$>$ where the goal is to determine the optimal selection of items to maximize the total $<$entity$>$ $<$delete$>$ value$<$/delete$>$ $<$mark$>$ weight$<$/mark$>$ $<$/entity$>$ within a given weight constraint.
 Table 17: Examples of generated data from automatic data creation.
@@
 Reordering passages gives 1.7% improvements, further emphasizing the importance of retrieval and careful pipeline design.
+Analysis on data scaling.
+We conduct ablations of different data scaling in Fava. Figure 7 shows the scaling results.
+Figure 7: Training Size